### v1

# Axis Tour: Word Tour Determines the Order of Axes in ICA-transformed Embeddings

Abstract
Word embedding is one of the most important components in natural language processing, but interpreting high-dimensional embeddings remains a challenging problem. To address this problem, Independent Component Analysis (ICA) is identified as an effective solution. ICA-transformed word embeddings reveal interpretable semantic axes; however, the order of these axes are arbitrary. In this study, we focus on this property and propose a novel method, Axis Tour, which optimizes the order of the axes. Inspired by Word Tour, a one-dimensional word embedding method, we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. Furthermore, we show through experiments on downstream tasks that Axis Tour constructs better low-dimensional embeddings compared to both PCA and ICA.
1 Introduction
Figure 1: Scatterplots of normalized ICA-transformed word embeddings whose axes are ordered by Axis Tour and Skewness Sort. In the upper part, Axis Tour is applied to 300-dimensional GloVe, with nine consecutive axes arranged counterclockwise. In the lower part, these nine axes are arranged clockwise in descending order of skewness. The embeddings are projected onto two dimensions along these axes. The top five embeddings on each axis are labeled by their words. Each word is assigned the color of the axis with the highest value. In both cases, words that cross the horizontal axes are not displayed.
Word embedding is one of the most important components in natural language processing, but interpreting high-dimensional embeddings remains a challenging problem. To address this problem, Independent Component Analysis (ICA) Hyvärinen and Oja (2000) is identified as an effective solution Mareček et al. (2020); Musil and Marecek (2022); Yamagiwa et al. (2023). ICA-transformed word embeddings reveal interpretable semantic axes; however, the order of these axes are arbitrary Hyvärinen et al. (2001b).
Inspired by a one-dimensional word embeddings, Word Tour Sato (2022), which leverages the Traveling Salesman Problem (TSP), we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes.
Fig. 1 shows two sets of two-dimensional projections of word embeddings: one set is ordered by Axis Tour and the other is sorted in descending order of skewness (Skewness Sort). In Axis Tour, the most significant words of the axes are far from the center, with the meanings of the axes changing continuously. Conversely, in Skewness Sort, the most significant words are nearer to the center, and axes with different meanings are placed adjacently. In fact, the average distance from the origin to these significant words in Fig. 1 is $0.76$ in Axis Tour, as opposed to $0.65$ in Skewness Sort.
We also assume that the consecutive axes in the Axis Tour embeddings can be considered as a subspace whose axes have similar meanings. Based on this, we project each subspace onto a single dimension to achieve dimensionality reduction. We show through experiments that Axis Tour yields better low-dimensional embeddings when compared to both PCA and ICA.
2 Related work
There are studies that transform embeddings by rotation Park et al. (2017) from Factor Analysis Crawford and Ferguson (1970); Browne (2001) or Principal Component Analysis (PCA) Musil (2019). Independent Component Analysis (ICA)  Hyvärinen and Oja (2000) has gained attention for its ability to reveal interpretable semantic axes in the transformed embeddings Mareček et al. (2020); Musil and Marecek (2022); Yamagiwa et al. (2023).
Relevant for our study is Topographic ICA (TICA) Kohonen (2001); Hyvärinen et al. (2001a). TICA relaxes the assumption of statistical independence and assumes energy correlations between adjacent axes. TICA then estimates the order of the axes. Unlike TICA, Axis Tour applies to ordinary ICA-transformed embeddings and uses the embeddings themselves to measure axis similarity.
3 Background
The pre-trained word embedding matrix is given by $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, where $\mathbf{X}$ is centered (i.e., the mean of each column is zero). Here, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the word embedding of the $i$-th word.
3.1 ICA-transformed word embeddings
ICA Hyvärinen and Oja (2000) finds the transformation matrix $\mathbf{B}\in\mathbb{R}^{d\times d}$ such that the columns of the matrix $\mathbf{S}\in\mathbb{R}^{n\times d}$, represented by the following equation, are as statistically independent as possible:
$\mathbf{S}=\mathbf{X}\mathbf{B},$ (1)
where $\mathbf{S}$ is assumed to be whitened (i.e., the variance of each column is $1$). The columns of $\mathbf{S}$ are called independent components [^1] Unless otherwise noted, flip the sign of each axis as needed so that the skewness is positive..
While the matrix $\mathbf{S}$ has interpretable semantic axes Mareček et al. (2020); Musil and Marecek (2022); Yamagiwa et al. (2023), the order of these axes are arbitrary Hyvärinen et al. (2001b).
3.2 Word Tour
Let $\mathcal{P}([n])$ be the set of all permutations of $[n]$, where $[n]=\{1,\cdots,n\}$. Word Tour Sato (2022) is a one-dimensional word embeddings method which solves the following Traveling Salesman Problem (TSP):
$\underset{\sigma\in\mathcal{P}([n])}{\operatorname{min}}\,\|\mathbf{x}_{\sigma_{1}}-\mathbf{x}_{\sigma_{n}}\|+\sum_{i=1}^{n-1}\|\mathbf{x}_{\sigma_{i}}-\mathbf{x}_{\sigma_{i+1}}\|.$ (2)
The resulting one-dimensional embeddings have similar meanings if they are close in order.
4 Axis Tour
This section explains Axis Tour and the dimensionality reduction method using Axis Tour. As mentioned in Section 3.2, $[d]=\{1,\cdots,d\}$ and $\mathcal{P}([d])$ is the set of all permutations of $[d]$.
4.1 Definition of axis embedding
We define axis embedding for use in Word Tour. The embedding represents the axis of the ICA-transformed embeddings $\mathbf{S}$.
In preparation, we define the matrix $\hat{\mathbf{S}}\in\mathbb{R}^{n\times d}$ as the normalization of the embeddings $\mathbf{S}$ with row vectors $\hat{\mathbf{s}}_{i}=\mathbf{s}_{i}/\|\mathbf{s}_{i}\|$. Here, the $i$-th word embedding of $\mathbf{S}$ and $\hat{\mathbf{S}}$ are denoted by $\mathbf{s}_{i},\hat{\mathbf{s}}_{i}\in\mathbb{R}^{d}$, respectively. We compare the elements of the $\ell$-th axis of $\hat{\mathbf{S}}$ and denote the index set of words corresponding to the top $k$ elements as $\text{Top}_{k}^{\ell}$.
We then define the $\ell$-th axis embedding $\mathbf{v}_{\ell}$ for $\mathbf{S}$ as follows:
$\mathbf{v}_{\ell}:=\frac{1}{k}\sum_{i\in\text{Top}_{k}^{\ell}}\hat{\mathbf{s}}_{i}\in\mathbb{R}^{d}.$ (3)
As we saw in Fig. 1, since the meaning of an axis can be interpreted from the top words, $\mathbf{v}_{\ell}$ can be considered as the embedding that represents the meaning of the $\ell$-th axis of $\mathbf{S}$.
4.2 Determining the order of axes
Axis Tour is a method that uses $\mathbf{v}_{\ell}$ in (3) to perform Word Tour and determines the order of axes in ICA-transformed word embeddings. In Axis Tour, the cost between the axis embeddings $\mathbf{v}_{\ell},\mathbf{v}_{m}$ used for TSP is defined by $1-\cos{(\mathbf{v}_{\ell},\mathbf{v}_{m})}$ instead of $\|\mathbf{v}_{\ell}-\mathbf{v}_{m}\|$. This approach then maximizes the sum of cosine similarities between adjacent axis embeddings. Therefore, the problem is formulated as follows [^2] Note that due to the cyclic nature of $\tau$, we ensure that $\cos(\mathbf{v}_{\tau_{1}},\mathbf{v}_{\tau_{d}})$ is the smallest of the cosine similarities.:
$\underset{\tau\in\mathcal{P}([d])}{\operatorname{max}}\,\cos(\mathbf{v}_{\tau_{1}},\mathbf{v}_{\tau_{d}})+\sum_{\ell=1}^{d-1}\cos(\mathbf{v}_{\tau_{\ell}},\mathbf{v}_{\tau_{\ell+1}}).$ (4)
The sum of cosine similarities between adjacent axis embeddings can be considered as a measure of the semantic continuity of the axes. Thus, Axis Tour determines the order of the axes by maximizing this measure.
4.3 Dimensionality reduction
[…]
First, we consider dimensionality reduction for $I_{r}$. To do this, we define a vector $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ of length $1$ for each $I_{r}$ as follows:
$f_{r}^{(\ell)}=\begin{cases}\gamma_{\ell}^{\alpha}/\sqrt{\sum_{m=a_{r}}^{b_{r}}\gamma_{m}^{2\alpha}}&\text{for }\ell\in I_{r}\\ 0&\text{otherwise},\end{cases}$ (5)
where $\alpha\in\mathbb{R}_{\geq 0}$. Then $\mathbf{T}\mathbf{f}_{r}\in\mathbb{R}^{n}$ can be considered as a projection of the subspace spanned by the axes of $\mathbf{T}$ corresponding to $I_{r}$ onto a one-dimensional space. Fig. 5 in Appendix A shows the projection for three consecutive axes.
Next, we define the matrix $\mathbf{F}:=[\mathbf{f}_{1},\cdots,\mathbf{f}_{p}]\in\mathbb{R}^{d\times p}$. Then $\mathbf{T}\mathbf{F}\in\mathbb{R}^{n\times p}$ represents the concatenated projections, that is, a dimensionality reduction of the $d$-dimensional embeddings $\mathbf{T}$ to $p$ dimensions. For more details, refer to Appendix A.
5 Experiments
23 24 25 26 27 28 29 30 31
serb russian czech germany france canada australia wiltshire liga
bosnian russia prague german french canadian australian shrewsbury relegated
croatia moscow poland berlin le ontario queensland lincolnshire fc
croatian sergei polish von paris quebec brisbane peterborough f.c.
serbian aleksandr warsaw cologne du saskatchewan perth croydon serie
101 102 103 104 105 106 107 108 109
pay land laws court lawsuits charges camp corpses remain
fees property regulations judge lawsuit alleged prison corpse remained
payments lands enacted appellate litigation prosecutors buchenwald exhumed stayed
payment estate law appeals suits indicted camps dismembered stubbornly
paid bergisches provisions supreme suit convicted inmates bodies stays
237 238 239 240 241 242 243 244 245
award film superhero album piano paintings manuscript language name
awards films marvel albums violin painting biographies languages names
awarded movie spin-off band cello art pages pashto surname
prize starring superheroes self-titled percussion sculpture book colloquial phrase
emmy directed characters ep orchestral watercolor handwritten dialect misspelling
Table 1: Semantic continuity of axes via Axis Tour for normalized ICA-transformed embeddings. We apply Axis Tour to 300-dimensional GloVe and extract nine consecutive axes, then display the top five words for each axis.
Figure 2: Histogram of the cosine similarity between consecutive axis embeddings $\cos(\mathbf{v}_{i},\mathbf{v}_{i+1})$. As an additional baseline, we sampled 300 random words, ordered them, and showed the histogram of adjacent cosine similarities. The dashed lines indicates the average similarity for each method. The distribution for Axis Tour shifts towards a more positive mean, distancing itself from 0. Conversely, other distributions roughly follow the normal distribution with means that are close to 0.
[…]
Figure 3: Relationship between the skewness and the average of consecutive two cosines (i.e., $(\cos(\mathbf{v}_{i-1},\mathbf{v}_{i})+\cos(\mathbf{v}_{i},\mathbf{v}_{i+1}))/2$) for (a) Axis Tour and (b) Skewness Sort. The left shows the skewness and the average of two cosines for both $y$-axes, and the right shows the scatter plot of these values. Spearman’s rank correlation is 0.43 for Axis Tour, while it is 0.04 for Skewness Sort.
Figure 4: Dimensionality reduction performance for embeddings. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks. For detailed experimental results, refer to Appendix B.
Similar to the Word Tour experiments, we used 300-dimensional GloVe  Pennington et al. (2014) with $n=40{,}000$, and the LKH solver [^5] The LKH solver is an implementation of Lin-Kernighan algorithm Lin and Kernighan (1973); Helsgaun (2000). Helsgaun (2018) as the TSP solver.
For ICA, we employed FastICA Hyvärinen (1999) from scikit-learn Pedregosa et al. (2011), setting iterations to $10{,}000$ and tolerance to $10^{-10}$, consistent with the setting in Yamagiwa et al. (2023). We computed the axis embeddings $\mathbf{v}_{\ell}$ in (3) with $k=100$. For baselines, we used whitened PCA-transformed embeddings [^6] Whitened ICA-transformed embeddings are obtained by applying an orthogonal matrix to these embeddings., along with two types of whitened ICA-transformed embeddings: Random Order, which randomizes the axes’ order in $\mathbf{S}$ [^7] In the embeddings matrix, we do not flip the sign of each axis to ensure positive skewness., and Skewness Sort, which arranges the axes of $\mathbf{S}$ in descending order of skewness. See Appendix C for additional experiments.
5.1 Qualitative observation
Table 1 shows examples of consecutive axes extracted using Axis Tour. These examples demonstrate the semantic similarity between adjacent axes and illustrate how the meanings of the axes change continuously. For example, in top row, the axis meaning transition from Eastern Europe to Germany and France, followed by Canada (which shares a connection with France) then to Australia (English-speaking regions), the regions in England, and finally to soccer (a popular sport in England), showcasing geographic and cultural continuity.
5.2 Cosine similarity between adjacent axis embeddings
Fig. 2 shows the histograms of $\cos(\mathbf{v}_{i},\mathbf{v}_{i+1})$ for Axis Tour and the baselines. In Axis Tour, the values of $\cos(\mathbf{v}_{i},\mathbf{v}_{i+1})$ are consistently higher, while this trend is not observed in the other baselines. This result is consistent with the formulation in (4).
For Axis Tour and Skewness Sort, Fig. 3 illustrates the relationship between the skewness and the average of the two adjacent cosine similarities [^8] Cosine similarity is not defined for the $i$-th axis itself.. In Fig. 3 for Axis Tour, it is interesting to note that these two values are correlated. However, no such trend is observed in Fig. 3 for Skewness Sort. This indicates that in Axis Tour, the axes with higher skewness are more likely to be ordered adjacent to the axes with higher similarity.
5.3 Dimensionality reduction: analogy, word similarity, and categorization tasks
[…]
Fig. 4 shows that the dimensionality reduction for Axis Tour is better than or equal to the baselines for most dimensions in each task. This result suggests that the dimensionality reduction for Axis Tour efficiently merges the axes with similar meanings. For more details, refer to Appendix B.
6 Conclusion
In this study, we propose a novel method, Axis Tour, which optimizes the order of axes in ICA-transformed word embeddings. We focus on the fact that the word embeddings reveal interpretable semantic axes while the order of these axes are arbitrary. Axis Tour aims to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. Furthermore, we show through experiments on downstream tasks that Axis Tour constructs better low-dimensional embeddings compared to both PCA and ICA.
Limitations
  • While the dimension reduction experiments showed the improvement of the downstream task performance for the Axis Tour embeddings, there are three aspects that could be further improved:
    1. 1.
      Dimension reduction is performed using the vector $\mathbf{f}_{r}$, but its definition (5) is empirical, and better vectors may be designed. In addition, nonlinear transformations beyond linear ones could be considered for dimension reduction. Details on the definition of $\mathbf{f}_{r}$ can be found in Appendix A.
    2. 2.
      The method in Section 4.3 simply divides $[d]$ into $p$ equal intervals to merge the axes. However, adaptively determining the division points could allow selecting more semantically coherent groups of axes.
    3. 3.
      To construct optimal low-dimensional vectors using ICA-transformed embeddings, applying clustering methods such as K-means to axis embeddings may improve performance. In this case, the overall optimized axis order may not be determined as in Axis Tour, but performing Axis Tour within each cluster and then concatenating these could determine an axis order depending on the number of clusters.
    However, this study focuses on a method for maximizing the semantic continuity of axes in ICA-transformed embeddings, leaving detailed investigation of the effective low-dimensional vector as future work.
  • In Axis Tour, while adjacent axes may have similar meanings, axes with similar meanings may not be in close order. This is due to the fact that in Word Tour, high-dimensional embeddings result in one-dimensional embeddings, and the meanings of words are similar when the word order is close, but semantically similar words are not always embedded close to each other.
  • As seen in Fig. 1, projecting multiple axes of ICA-transformed embeddings into two dimensions can effectively represent the shape of the embeddings. However, as the number of axes increases, the angles between the axes become small, resulting in crowded axes. This can cause problems such as the top words of the axes being closer to the origin, which can be difficult to interpret.
  • In Axis Tour, the dimension of the ICA-transformed embeddings corresponds to the number of cities in TSP. Therefore, as the dimension of the embeddings increases, the computation time for Axis Tour becomes longer. Note that for the 300-dimensional GloVe used in this study, the computation time for Axis Tour is about one second. For reference, Word Tour with $n=40{,}000$ is known to take several hours [^11] https://github.com/joisino/wordtour..
Ethics Statement
This study complies with the ACL Ethics Policy.
Acknowledgements
We would like to thank Momose Oyama for the discussion. This work was supported by JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2123. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3.
Appendix A Dimensionality reduction
Figure 5: Projection of the subspace spanned by three consecutive axes in Fig. 1 into a one-dimensional space. Each word is assigned the color of the axis with the highest value. The projection direction is in the direction representing the subspace. For visualization, we randomly sampled 10,000 words, excluding the top 5 words on each axis.
This section details and supplements Section 4.3.
A.1 Definition of $\mathbf{f}_{r}$
Here we first discuss the definition of $\mathbf{f}_{r}$ in terms of skewness and then explain the normalization.
Skewness as weight.
In (3), a vector $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for dimensionality reduction is defined by the skewness $\gamma_{\ell}\in\mathbb{R}_{\geq 0}$ of the $\ell$-th axis of $\mathbf{T}$. In particular, $f_{r}^{(\ell)}$, which corresponds to the weight of the $\ell$-th axis in the projection, is proportional to $\gamma_{\ell}^{\alpha}(\alpha\in\mathbb{R}_{\geq 0})$. This is based on the assumption that the axis becomes more meaningful as the skewness increases [^12] While not specific to ICA-transformed embeddings, it is known from a study of sparse coding for language models that skewness correlates with interpretability Cunningham et al. (2023).. Higher-order statistics such as skewness are known to be sensitive to outliers Hyvärinen et al. (2001b), so we mitigate this effect by raising $\gamma_{\ell}$ to the power of $\alpha$. Since skewness is a third-order moment, we treat $\alpha=1/3$ as the default value. $\alpha=0$ is a uniform weight, while $\alpha=1$ is the original $\gamma_{\ell}$. Next, we will explain why $\mathbf{f}_{r}$ is normalized.
Reason for Normalizing $\mathbf{f}_{r}$.
$\mathbf{f}_{r}$ is normalized (i.e., $\|\mathbf{f}_{r}\|=1$) to equalize the scale of the one-dimensional projections for each subspace. Thus, when $p=d$, $\mathbf{T}\mathbf{F}=\mathbf{T}$ holds.
For $p=d$, $I_{r}=\{r\}$, and from (3), $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}$ satisfies:
$f_{r}^{(\ell)}=\delta_{\ell r}=\begin{cases}1&\ell=r\\ 0&\ell\neq r\end{cases}.$ (6)
Here, $\delta_{\ell r}$ is the Dirac delta function. From (6) the matrix $\mathbf{F}=[\mathbf{f}_{1},\cdots,\mathbf{f}_{d}]\in\mathbb{R}^{d\times d}$ corresponds to the $d$-dimensional identity matrix $\mathbf{I}\in\mathbb{R}^{d\times d}$, so $\mathbf{T}\mathbf{F}=\mathbf{T}\mathbf{I}=\mathbf{T}$.
A.2 Projection from subspace to one-dimensional space
This section explains the projection from a subspace to a one-dimensional space using a specific example. Consider the subspace spanned by three consecutive axes (89, 90, 91) from Fig. 1. Fig. 5 shows the projection of this subspace using $\mathbf{f}_{r}$. The projection direction is in the direction representing the subspace, and the top words of each axis are projected close together.
Appendix B Details of experiment in Section 5
$p=5$ $p=20$ $p=100$ $p=300$ Tasks PCA Rand. Skew. Tour. PCA Rand. Skew. Tour. PCA Rand. Skew. Tour. All Analogy capital-common-countries 0.00 0.00 0.00 0.00 0.37 0.00 0.06 0.11 0.95 0.56 0.85 0.87 0.95 capital-world 0.00 0.00 0.00 0.00 0.26 0.02 0.05 0.11 0.90 0.73 0.80 0.82 0.95 city-in-state 0.00 0.00 0.00 0.00 0.03 0.00 0.02 0.04 0.42 0.28 0.24 0.40 0.67 currency 0.00 0.00 0.00 0.00 0.02 0.00 0.00 0.02 0.09 0.08 0.08 0.10 0.12 family 0.01 0.00 0.00 0.00 0.40 0.02 0.08 0.22 0.78 0.68 0.80 0.75 0.88 gram1-adjective-to-adverb 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.01 0.08 0.08 0.14 0.09 0.21 gram2-opposite 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.14 0.12 0.18 0.14 0.26 gram3-comparative 0.00 0.00 0.00 0.00 0.05 0.01 0.04 0.11 0.62 0.46 0.58 0.66 0.88 gram4-superlative 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.01 0.31 0.54 0.23 0.31 0.69 gram5-present-participle 0.00 0.00 0.00 0.00 0.03 0.00 0.03 0.07 0.44 0.30 0.59 0.58 0.69 gram6-nationality-adjective 0.00 0.00 0.00 0.00 0.57 0.07 0.22 0.43 0.91 0.88 0.91 0.88 0.93 gram7-past-tense 0.00 0.00 0.00 0.00 0.05 0.02 0.04 0.06 0.45 0.36 0.47 0.51 0.60 gram8-plural 0.00 0.00 0.00 0.00 0.11 0.01 0.04 0.07 0.73 0.40 0.59 0.56 0.76 gram9-plural-verbs 0.00 0.00 0.00 0.00 0.06 0.01 0.02 0.06 0.39 0.27 0.29 0.53 0.58 jj_jjr 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.03 0.35 0.23 0.29 0.43 0.66 jj_jjs 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.01 0.21 0.36 0.14 0.20 0.51 jjr_jj 0.00 0.00 0.00 0.00 0.01 0.00 0.03 0.02 0.33 0.27 0.32 0.33 0.54 jjr_jjs 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.01 0.24 0.37 0.15 0.20 0.55 jjs_jj 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.01 0.21 0.13 0.18 0.19 0.48 jjs_jjr 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.01 0.29 0.16 0.25 0.33 0.63 nn_nnpos 0.00 0.00 0.00 0.00 0.05 0.01 0.01 0.04 0.35 0.20 0.29 0.28 0.42 nn_nns 0.00 0.00 0.00 0.00 0.04 0.00 0.04 0.06 0.55 0.33 0.49 0.51 0.74 nnpos_nn 0.00 0.00 0.00 0.00 0.03 0.00 0.02 0.06 0.40 0.20 0.34 0.31 0.45 nns_nn 0.00 0.00 0.00 0.00 0.05 0.00 0.04 0.06 0.48 0.30 0.43 0.44 0.64 vb_vbd 0.00 0.00 0.00 0.00 0.11 0.03 0.09 0.07 0.45 0.36 0.40 0.54 0.58 vb_vbz 0.00 0.00 0.00 0.00 0.09 0.02 0.05 0.08 0.58 0.33 0.50 0.68 0.76 vbd_vb 0.00 0.00 0.00 0.00 0.09 0.04 0.12 0.08 0.46 0.36 0.56 0.56 0.69 vbd_vbz 0.00 0.00 0.00 0.00 0.08 0.01 0.04 0.05 0.48 0.26 0.38 0.54 0.63 vbz_vb 0.00 0.00 0.00 0.00 0.06 0.02 0.06 0.11 0.65 0.43 0.70 0.70 0.82 vbz_vbd 0.00 0.00 0.00 0.00 0.05 0.01 0.03 0.05 0.33 0.33 0.42 0.44 0.55 Average 0.00 0.00 0.00 0.00 0.09 0.01 0.04 0.07 0.45 0.34 0.42 0.46 0.63 Similarity MEN 0.16 0.19 0.11 0.35 0.32 0.33 0.29 0.51 0.66 0.56 0.63 0.66 0.75 MTurk 0.17 0.08 0.12 0.32 0.38 0.32 0.30 0.52 0.57 0.53 0.57 0.61 0.64 RG65 0.31 0.09 0.05 0.29 0.36 0.42 0.28 0.50 0.68 0.66 0.59 0.63 0.78 RW 0.09 0.10 0.07 0.13 0.14 0.16 0.10 0.25 0.24 0.32 0.28 0.30 0.34 SimLex999 0.01 0.13 0.04 0.07 0.11 0.21 0.08 0.21 0.27 0.37 0.28 0.31 0.40 WS353 0.12 0.16 0.02 0.31 0.15 0.28 0.18 0.44 0.47 0.40 0.43 0.52 0.57 WS353R 0.12 0.14 0.01 0.17 0.15 0.16 0.15 0.35 0.40 0.28 0.35 0.44 0.51 WS353S 0.18 0.21 0.06 0.45 0.21 0.35 0.26 0.55 0.57 0.51 0.58 0.62 0.69 Average 0.15 0.14 0.06 0.26 0.23 0.28 0.20 0.42 0.48 0.45 0.46 0.51 0.59 Categorization AP 0.33 0.22 0.22 0.27 0.36 0.26 0.28 0.40 0.51 0.45 0.54 0.56 0.66 BLESS 0.31 0.28 0.27 0.36 0.42 0.36 0.35 0.51 0.73 0.68 0.69 0.76 0.79 Battig 0.18 0.10 0.12 0.15 0.24 0.14 0.16 0.22 0.37 0.29 0.35 0.34 0.42 ESSLI_1a 0.50 0.41 0.45 0.64 0.61 0.57 0.48 0.77 0.73 0.59 0.73 0.75 0.70 ESSLI_2b 0.47 0.62 0.45 0.53 0.73 0.68 0.55 0.68 0.75 0.75 0.70 0.75 0.78 ESSLI_2c 0.44 0.44 0.42 0.44 0.53 0.60 0.49 0.56 0.58 0.53 0.60 0.60 0.58 Average 0.37 0.35 0.32 0.40 0.48 0.44 0.38 0.52 0.61 0.55 0.60 0.63 0.65
Table 2: Dimensionality reduction performance for $p$-dimensional embeddings. Rand. stands for Random Order, Skew. for Skewness Sort, and Tour. for Axis Tour. The values in the table correspond to top 1 accuracy for analogy tasks, Spearman’s rank correlation for word similarity tasks, and purity for categorization tasks. Note that at $p=300$, all embeddings give the same results.
Figure 6: Dimensionality reduction performance for the PCA-transformed embeddings and the Axis Tour embeddings whose $\alpha=0,1/3,1/2,1$. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks.
The embedding of the word $w_{i}$ is denoted by $\mathbf{w}_{i}\in\mathbb{R}^{d}$.
Analogy task.
We used the Google Analogy Test Set Mikolov et al. (2013a), which contains 14 types of analogy tasks, and the Microsoft Research Syntactic Analogies Dataset Mikolov et al. (2013b), which contains 16 types of analogy tasks. In the analogy tasks, the quality of the embeddings is evaluated by inferring $w_{4}$ to which $w_{3}$ corresponds if $w_{1}$ corresponds to $w_{2}$. We compute the vector $\mathbf{w}_{2}-\mathbf{w}_{1}+\mathbf{w}_{3}$ and see if the closest embedding is $\mathbf{w}_{4}$ (top1 accuracy).
Word similarity task.
We used MEN Bruni et al. (2014), MTurk Radinsky et al. (2011), RG65 Rubenstein and Goodenough (1965), RW Luong et al. (2013), SimLex999 Hill et al. (2015), WS353 Finkelstein et al. (2002), WS353R (WS353 Relatedness), and WS353S (WS353 Similarity). In the word similarity tasks, the quality of the embeddings is evaluated by measuring the cosine similarity of the word embeddings and comparing it to the human-rated similarity scores. As the evaluation metric, we used Spearman’s rank correlation coefficient between the human ratings and the cosine similarity.
Categorization task.
We used AP Almuhareb and Poesio (2005), BLESS Baroni and Lenci (2011), Battig Battig and Montague (1969), ESSLLI_1a Baroni et al. (2008), ESSLLI_2b Baroni et al. (2008), and ESSLLI_2c Baroni et al. (2008). In the categorization tasks, the quality of the embeddings is evaluated by clustering them in the setting where each word is assigned a class label. As the evaluation metric, we used Purity, which shows the proportion of the most frequent class in the clusters. As clustering methods, we used Hierarchical Clustering with five settings [^13] By default, Word Embedding Benchmark uses the following affinity and linkage pairs for hierarchical clustering: (affinity, linkage) = (euclidean, ward), (euclidean, average), (euclidean, complete), (cosine, average), (cosine, complete). and K-means [^14] We used the same seed for all experiments., and then selected the one that gave the highest purity.
Results.
Table 2 shows detailed experimental results of PCA, Random Order, Skewness Sort, and Axis Tour at $p=5,20,100,300$ for each task. As already seen, Fig. 4 in Section 5 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the embeddings.
The Axis Tour embeddings showed superior performance in the word similarity tasks and the categorization tasks for almost all dimensions compared to other methods. In the analogy tasks, the Axis Tour embeddings achieved performance comparable to PCA and better than Random Order and Skewness Sort in most dimensions.
Note that the experimental results are the same for all embeddings for $p=300$. First, as we saw in Appendix A.1, when $p=d\,(=300$), the matrix $\mathbf{TF}$ (i.e., the projected $p$-dimensional embeddings) is equal to the matrix $\mathbf{T}$ (i.e., the $d$-dimensional Axis Tour embeddings). Then, by definition, Axis Tour, Random Order, and Skewness Sort are the embeddings obtained by reordering the axes of the ICA-transformed embeddings and flipping their signs as needed. Thus, these three can be seen as the embeddings obtained by applying an orthogonal matrix to the ICA-transformed embeddings. Since the ICA-transformed embeddings are derived from the PCA-transformed embeddings by applying an orthogonal matrix [^15] Refer to the previous work for the relationship between PCA and ICA Yamagiwa et al. (2023)., cosine similarity and Euclidean distance remain unchanged for PCA, Random Order, Skewness Sort, and Axis Tour, leading to identical results in downstream tasks.
Appendix C Additional experiments
C.1 Comparison of $\alpha$
Fig. 6 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the PCA-transformed embeddings and the Axis Tour embeddings whose $\alpha=0,1/3,1/2,1$.
From Fig. 6, we can see that the performance of the Axis Tour embeddings changes for each task, depending on $\alpha$. For example, when comparing across all $\alpha$, while $\alpha=1$ shows good performance on analogy tasks and poor performance on word similarity tasks, $\alpha=0$ shows the opposite. These results indicate that the quality of low-dimensional embeddings by the Axis Tour embeddings depends on the vector $\mathbf{f}_{r}$ for projection. However, the overall changes for each alpha are not as large, and in all tasks the performance is comparable to or better than that of the PCA-transformed embeddings, indicating the ability to construct high-quality low-dimensional embeddings.