### v3

# Axis Tour: Word Tour Determines the Order of Axes in ICA-transformed Embeddings

Abstract
Word embedding is one of the most important components in natural language processing, but interpreting high-dimensional embeddings remains a challenging problem. To address this problem, Independent Component Analysis (ICA) is identified as an effective solution. ICA-transformed word embeddings reveal interpretable semantic axes; however, the order of these axes are arbitrary. In this study, we focus on this property and propose a novel method, Axis Tour, which optimizes the order of the axes. Inspired by Word Tour, a one-dimensional word embedding method, we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. Furthermore, we show through experiments on downstream tasks that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
1 Introduction
Figure 1: Scatterplots of normalized ICA-transformed word embeddings whose axes are ordered by Axis Tour and Skewness Sort. In the upper part, Axis Tour is applied to 300-dimensional GloVe, with nine consecutive axes arranged counterclockwise. In the lower part, these nine axes are rearranged clockwise in descending order of skewness. The embeddings are projected onto two dimensions along these axes. The top five embeddings on each axis are labeled by their words. Each word is assigned the color of the axis with the highest value. In both cases, words that cross the horizontal axes are not displayed. Refer to Appendix A for more details.
Embedding is an important tool in natural language processing, but interpreting high-dimensional embeddings is challenging. To address this, Independent Component Analysis (ICA) Hyvärinen and Oja (2000) offers an effective solution Mareček et al. (2020); Musil and Mareček (2024); Yamagiwa et al. (2023). ICA-transformed embeddings reveal interpretable semantic axes; however, the order of these axes is arbitrary Hyvärinen et al. (2001b). In this study, inspired by a one-dimensional word embedding method, Word Tour Sato (2022), which leverages the Traveling Salesman Problem (TSP), we aim to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes.
Figure 1 shows two sets of two-dimensional projections of word embeddings: one is ordered by Axis Tour (our proposal), and the other is sorted in descending order of skewness (Skewness Sort). In Axis Tour, the top words of the axes are positioned farther from the center, with the meanings of the axes changing continuously. Conversely, in Skewness Sort, the top words are closer to the center, and the axes with different meanings are placed adjacently. In fact, the average distance from the origin to the top words in Fig. 1 is $0.76$ in Axis Tour, compared to $0.61$ in Skewness Sort.
We also assume that the consecutive axes in the Axis Tour embeddings can be considered a subspace with axes that have similar meanings. Based on this idea, we project each subspace onto a single dimension for dimensionality reduction. We show through experiments that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
2 Related work
Some studies transform embeddings by rotation Park et al. (2017) from Factor Analysis Crawford and Ferguson (1970); Browne (2001) or Principal Component Analysis (PCA) Musil (2019). Independent Component Analysis (ICA) Hyvärinen and Oja (2000) has gained attention for its ability to reveal interpretable semantic axes in the transformed embeddings Mareček et al. (2020); Musil and Mareček (2024); Yamagiwa et al. (2023).
Research on interpreting embeddings by focusing on axes representing opposing concepts (e.g., cold vs. hot, soft vs. hard) is also actively pursued. Approaches such as SemAxis An et al. (2018), POLAR Mathew et al. (2020), and FrameAxis Kwak et al. (2021) deal with static embeddings, while BiImp Senel et al. (2022) and SensePOLAR Engler et al. (2022) deal with dynamic embeddings. In particular, Section 5.3 provides a comparison of the Axis Tour embeddings and those from POLAR.
Relevant to our study is Topographic ICA (TICA) Kohonen (2001); Hyvärinen et al. (2001a). TICA relaxes the assumption of statistical independence and assumes higher-order correlations between adjacent axes. then estimates the order of the axes. Unlike TICA, Axis Tour is applied to ordinary ICA-transformed embeddings and uses the embeddings themselves to measure axis similarity. For more details on TICA, refer to Appendix F.
3 Background
The pre-trained word embedding matrix is given by $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, where $\mathbf{X}$ is centered (i.e., the mean of each column is zero). Here, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the word embedding of the $i$-th word.
3.1 ICA-transformed word embeddings
ICA Hyvärinen and Oja (2000) finds the transformation matrix $\mathbf{B}\in\mathbb{R}^{d\times d}$ such that the columns of the matrix $\mathbf{S}\in\mathbb{R}^{n\times d}$, represented by the following equation, are as independent as possible:
$\mathbf{S}=\mathbf{X}\mathbf{B},$ (1)
where $\mathbf{S}$ is whitened (i.e., the variances of the columns are $1$ and their correlations are all 00). The columns of $\mathbf{S}$ are called independent components [^1] Unless otherwise noted, flip the sign of each axis as needed so that the skewness is positive.. While $\mathbf{S}$ has interpretable semantic axes Mareček et al. (2020); Musil and Mareček (2024); Yamagiwa et al. (2023), the order of these axes are arbitrary Hyvärinen et al. (2001b).
3.2 Word Tour
Let $\mathcal{P}([n])$ be the set of all permutations of $[n]$, where $[n]=\{1,\ldots,n\}$. Word Tour Sato (2022) is a one-dimensional word embedding method that solves the following Traveling Salesman Problem (TSP):
$\underset{\tau\in\mathcal{P}([n])}{\operatorname{min}}\,\|\mathbf{x}_{\tau_{1}}-\mathbf{x}_{\tau_{n}}\|+\sum_{i=1}^{n-1}\|\mathbf{x}_{\tau_{i}}-\mathbf{x}_{\tau_{i+1}}\|.$ (2)
The resulting one-dimensional embeddings have similar meanings when they are close in order.
4 Axis Tour
This section explains Axis Tour and the dimensionality reduction method using Axis Tour. As mentioned in Section 3.2, $[d]=\{1,\ldots,d\}$ and $\mathcal{P}([d])$ is the set of all permutations of $[d]$.
4.1 Definition of axis embedding
We define axis embedding for use in Word Tour. The embedding represents the meaning of the axis of the ICA-transformed embeddings $\mathbf{S}$.
In preparation, we define the normalized ICA-transformed embeddings $\hat{\mathbf{S}}\in\mathbb{R}^{n\times d}$ as the normalization of the embeddings $\mathbf{S}$, where the row vectors are given by $\hat{\mathbf{s}}_{i}=\mathbf{s}_{i}/\|\mathbf{s}_{i}\|$. Here, the $i$-th word embeddings of $\mathbf{S}$ and $\hat{\mathbf{S}}$ are denoted by $\mathbf{s}_{i}$ and $\hat{\mathbf{s}}_{i}\in\mathbb{R}^{d}$, respectively. We compare the elements of the $\ell$-th axis of $\hat{\mathbf{S}}$ and denote the index set of words corresponding to the top $k$ elements as $\text{Top}_{k}^{\ell}$. We then define the $\ell$-th axis embedding $\mathbf{v}_{\ell}$ for $\mathbf{S}$ as follows:
$\mathbf{v}_{\ell}:=\frac{1}{k}\sum_{i\in\text{Top}_{k}^{\ell}}\hat{\mathbf{s}}_{i}\in\mathbb{R}^{d}.$ (3)
As we saw in Fig. 1, since the meaning of an axis can be interpreted from the top words, $\mathbf{v}_{\ell}$ can be considered the embedding that represents the meaning of the $\ell$-th axis of $\mathbf{S}$.
23 24 25 26 27 28 29 30 31
serb russian czech germany france canada australia wiltshire liga
bosnian russia prague german french canadian australian shrewsbury relegated
croatia moscow poland berlin le ontario queensland lincolnshire fc
croatian sergei polish von paris quebec brisbane peterborough f.c.
serbian aleksandr warsaw cologne du saskatchewan perth croydon serie
101 102 103 104 105 106 107 108 109
pay land laws court lawsuits charges camp corpses remain
fees property regulations judge lawsuit alleged prison corpse remained
payments lands enacted appellate litigation prosecutors buchenwald exhumed stayed
payment estate law appeals suits indicted camps dismembered stubbornly
paid bergisches provisions supreme suit convicted inmates bodies stays
237 238 239 240 241 242 243 244 245
award film superhero album piano paintings manuscript language name
awards films marvel albums violin painting biographies languages names
awarded movie spin-off band cello art pages pashto surname
prize starring superheroes self-titled percussion sculpture book colloquial phrase
emmy directed characters ep orchestral watercolor handwritten dialect misspelling
Table 1: Semantic continuity of axes by Axis Tour for normalized ICA-transformed embeddings. We apply Axis Tour to 300-dimensional GloVe and show the top five words for each axis. See Appendix E.1 for all axes results.
4.2 Determining the order of axes
Axis Tour is a method that uses $\mathbf{v}_{\ell}$ in (3) to perform Word Tour and determines the order of axes in ICA-transformed word embeddings. In Axis Tour, the cost between the axis embeddings $\mathbf{v}_{\ell}$ and $\mathbf{v}_{m}$ for the TSP is defined by $1-\cos{(\mathbf{v}_{\ell},\mathbf{v}_{m})}$ instead of $\|\mathbf{v}_{\ell}-\mathbf{v}_{m}\|$. This approach then maximizes the sum of cosine similarities between adjacent axis embeddings. Therefore, the problem is formulated as follows [^2] Note that due to the cyclic nature of $\tau$, we set $\tau_{1}$ such that $\cos(\mathbf{v}_{\tau_{1}},\mathbf{v}_{\tau_{d}})$ is the smallest of the cosine similarities.:
$\underset{\tau\in\mathcal{P}([d])}{\operatorname{max}}\,\cos(\mathbf{v}_{\tau_{1}},\mathbf{v}_{\tau_{d}})+\sum_{\ell=1}^{d-1}\cos(\mathbf{v}_{\tau_{\ell}},\mathbf{v}_{\tau_{\ell+1}}).$ (4)
The sum of cosine similarities between adjacent axis embeddings can be considered as a metric of the semantic continuity of the axes. Thus, Axis Tour determines the order of the axes by maximizing this metric.
4.3 Dimensionality reduction
[…]
First, we consider reducing the dimensionality of $\mathbf{T}$ along $I_{r}$, $r=1,\ldots,p$. To do this, we define a unit vector $\mathbf{f}_{r}:=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for each $I_{r}$ as follows:
$f_{r}^{(\ell)}=\begin{cases}\gamma_{\ell}^{\alpha}/\sqrt{\sum_{m=a_{r}}^{b_{r}}\gamma_{m}^{2\alpha}}&\text{for }\ell\in I_{r}\\ 0&\text{otherwise},\end{cases}$ (5)
where $\alpha\in\mathbb{R}_{\geq 0}$. Then $\mathbf{T}\mathbf{f}_{r}\in\mathbb{R}^{n}$ can be considered as a projection of the subspace spanned by the axes of $\mathbf{T}$ corresponding to $I_{r}$ onto a one-dimensional space. Fig. 7 in Appendix B shows the projection for three consecutive axes.
Next, we define the matrix $\mathbf{F}:=[\mathbf{f}_{1},\ldots,\mathbf{f}_{p}]\in\mathbb{R}^{d\times p}$. Then $\mathbf{T}\mathbf{F}\in\mathbb{R}^{n\times p}$ represents the concatenated projections, that is, a dimensionality reduction of the $d$-dimensional embeddings $\mathbf{T}$ to $p$ dimensions. For more details, refer to Appendix B.
5 Experiments
Similar to the Word Tour experiments, we used 300-dimensional GloVe  Pennington et al. (2014) with $n=400{,}000$, and the LKH solver [^5] The LKH solver is an implementation of Lin-Kernighan algorithm Lin and Kernighan (1973); Helsgaun (2000). Helsgaun (2018) for the optimization of the TSP.
For ICA, we used FastICA Hyvärinen (1999) from scikit-learn Pedregosa et al. (2011), setting the iterations to $10{,}000$ and the tolerance to $10^{-10}$, consistent with Yamagiwa et al. (2023). We computed the axis embeddings [^6] See Appendix E.3 for a discussion of the choice of $k$. $\mathbf{v}_{\ell}$ in (3) with $k=100$. For baselines, we used whitened PCA-transformed embeddings [^7] Whitened ICA-transformed embeddings are obtained by applying an orthogonal matrix to these embeddings., along with two types of whitened ICA-transformed embeddings [^8] See Appendix F for comparisons of Axis Tour and TICA.: Random Order, which randomly flips the sign of the axes in $\mathbf{S}$ and randomizes the order of the axes, and Skewness Sort, which sorts the axes of $\mathbf{S}$ in descending order of skewness. See Appendix E for additional experiments, including those of other embeddings.
5.1 Qualitative observation of semantic continuity in axis order
Table 1 presents an illustrative example of consecutive axes of the Axis Tour embeddings, where the three rows correspond to the meanings of countries, law, and art, respectively. We observe that the meanings of the axes change continuously. For instance, in the top row, the axis meaning shifts from Eastern Europe to Germany and France, followed by Canada (which shares a connection with France), then to Australia (English-speaking regions), the regions in England, and finally to soccer (a popular sport in England), demonstrating geographic and cultural continuity.
5.2 Quantitative evaluation of semantic continuity in axis order
This section quantitatively evaluates the semantic continuity of the axes in Axis Tour embeddings.
5.2.1 Evaluation by cosine similarity
First, we evaluate the two orderings of axes shown in Fig. 1. The semantic continuity of these axes is assessed by calculating the average cosine similarity between adjacent axis embeddings. For Axis Tour, the average cosine similarity is 0.269, but it decreases to 0.185 when these axes are rearranged by skewness, confirming the higher semantic continuity of the axes in Axis Tour.
Next, we consider the whole $d\,(=300)$ consecutive axes. Figure 2 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ for Axis Tour and the baselines. In Axis Tour, the values of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ are consistently higher, while this trend is not observed in the other baselines. The average cosine similarity is 0.244 for Axis Tour, while it is only 0.017 for Skewness Sort. This result is consistent with the formulation in (4).
5.2.2 Evaluation by GPT models
We also evaluate the semantic continuity of axes for Axis Tour and Skewness Sort using the OpenAI API. By focusing on the common axes in each embedding, we ask the model to determine, based on the top 10 words, whether the next axis of Axis Tour or that of Skewness Sort is more semantically related. The number of queries corresponds to $d\,(=300)$, and we use four GPT models: GPT-3.5 Turbo, GPT-4 Turbo, GPT-4o, and GPT-4o mini (see Appendix G for model versions and prompts).
As shown in Fig. 3, Axis Tour has a greater number of related axes compared to Skewness Sort for each model, implying more continuous changes in axis meanings. The smallest difference was observed with GPT-3.5 Turbo, the least performant model. For the other models (i.e., the GPT-4 models), the difference was at least four times larger.
Figure 2: Histogram of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$. As an additional baseline, we sampled 300 random words from the Random Order embeddings and arranged them in random order. The dashed lines represent the average similarity for each method. The distribution for Axis Tour shifts towards a more positive mean, while the others roughly follow a normal distribution with means close to 0. For more details, refer to Appendix C.
Figure 3: Comparison of the number of related axes in the GPT models. In each model, Axis Tour exhibits a greater number of related axes than Skewness Sort.
Figure 4: The performance of dimensionality reduction for embeddings. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks. See Appendix D for detailed experimental results.
(a) Axis Tour
(b) Skewness Sort
Figure 5: Relationship between the skewness $\gamma_{\ell}$ and the average of two consecutive cosines $(\cos(\mathbf{v}_{\ell-1},\mathbf{v}_{\ell})+\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1}))/2$ for all the axes $\ell=1,\ldots,d$ in (a) Axis Tour and (b) Skewness Sort. The left plot shows the skewness and the average of two cosines on both $y$-axes, while the right plot shows the scatter plot of these values. Spearman’s rank correlation is 0.43 for Axis Tour, while it is 0.04 for Skewness Sort.
5.3 Dimensionality reduction: analogy, word similarity, and categorization tasks
Using Word Embedding Benchmark Jastrzebski et al. (2017) [^9] https://github.com/kudkudak/word-embeddings-benchmarks., we evaluate the performance of dimensionality reduction in analogy, word similarity, and categorization tasks. PCA selects the axes in descending order of eigenvalue. Random Order and Skewness Sort select the axes sequentially from first to last. Axis Tour adopts the dimensionality reduction [^10] Fig. 8 in Appendix E.2 shows the results for different $\alpha$. in Section 4.3 with $\alpha=1/3$.
We consider the original GloVe embeddings $\mathbf{X}$ as well as the embeddings obtained by applying POLAR to $\mathbf{X}$. POLAR is a method that uses pairs of words with opposite meanings and finds the axes where these words are positioned at the opposite ends. There are three methods for selecting the axes: Random Selection, Variance Maximization, and Orthogonality Maximization. We used the publicly available code for our experiments. For both the original and POLAR-applied embeddings, axes were selected sequentially from first to last.
Figure 4 shows that the dimensionality reduction with the ordering in Axis Tour is better than or comparable to the baselines for most dimensionalities in each task. This result suggests that Axis Tour efficiently merges axes with similar meanings. For more details, refer to Appendix D.
6 Discussion
We confirmed through both qualitative and quantitative experiments that the axis order determined by Axis Tour exhibits high semantic continuity, and its effectiveness was also validated in the downstream task of dimensionality reduction.
There are two advantages to ordering the axes of ICA-transformed embeddings using Axis Tour. First, as shown in Fig. 1, when projecting the embedding space, the scatterplot becomes easier to interpret and more visually accessible because higher-ranking words on each axis are farther from the origin. Second, although the Axis Tour embeddings $\mathbf{T}$ are the same size as the ICA-transformed embeddings $\mathbf{S}$, the axis order in Axis Tour preserves information about the similarities between the axes.
An interesting relationship between semantic continuity and skewness is shown in Fig. 5, where the semantic continuity of axis $\ell$ is measured by the average of two cosines. In Axis Tour, there is a high correlation between skewness and semantic continuity, while in Skewness Sort, they are almost uncorrelated. A large skewness indicates that an axis has a distinctive meaning, and in Axis Tour, this seems to contribute to semantic continuity.
7 Conclusion
In this study, we proposed a novel method, Axis Tour, which optimizes the order of axes in ICA-transformed word embeddings. We focused on the fact that the word embeddings reveal interpretable semantic axes while the order of these axes is arbitrary. Axis Tour aims to improve the clarity of the word embedding space by maximizing the semantic continuity of the axes. Additionally, we demonstrated through experiments on downstream tasks that Axis Tour yields better or comparable low-dimensional embeddings compared to both PCA and ICA.
Limitations
  • While the dimension reduction experiments showed the improvement of the downstream task performance for the Axis Tour embeddings, there are three aspects that could be further improved:
    1. 1.
      Dimension reduction is performed using the vector $\mathbf{f}_{r}$, but its definition (5) is empirical, and better vectors may be designed. In addition, nonlinear transformations beyond linear ones could be considered for dimension reduction. Details on the definition of $\mathbf{f}_{r}$ can be found in Appendix B.
    2. 2.
      The method in Section 4.3 simply divides $[d]$ into $p$ equal intervals to merge the axes. However, adaptively determining the division points could allow selecting more semantically coherent groups of axes.
    3. 3.
      To construct optimal low-dimensional vectors using ICA-transformed embeddings, applying clustering methods such as $K$-means to axis embeddings may improve performance. In this case, the overall optimized axis order may not be determined as in Axis Tour, but performing Axis Tour within each cluster and then concatenating these could determine an axis order depending on the number of clusters.
    However, this study focuses on a method for maximizing the semantic continuity of axes in ICA-transformed embeddings, leaving detailed investigation of the effective low-dimensional vector as future work.
  • In Axis Tour, while adjacent axes may have similar meanings, axes with similar meanings may not be in close order. This is due to the fact that in Word Tour, high-dimensional embeddings result in one-dimensional embeddings, and the meanings of words are similar when the word order is close, but semantically similar words are not always embedded close to each other.
  • As seen in Fig. 1, projecting multiple axes of ICA-transformed embeddings into two dimensions can effectively represent the shape of the embeddings. However, as the number of axes increases, the angles between the axes become small, resulting in crowded axes. This can cause problems such as the top words of the axes being closer to the origin, which can be difficult to interpret.
  • In Axis Tour, the dimension of the ICA-transformed embeddings corresponds to the number of cities in TSP. Therefore, as the dimension of the embeddings increases, the computation time for Axis Tour becomes longer. Note that for the 300-dimensional GloVe used in this study, the computation time for Axis Tour is about one second. For reference, Word Tour with $n=40{,}000$ is known to take several hours [^11] https://github.com/joisino/wordtour..
Ethics Statement
A potential risk of this method is that we interpret the meanings of the axes of the ICA-transformed embeddings by the top words of each axis. If the embeddings contain personal information, such as email addresses or phone numbers, and these are contained in the top words, this can be problematic. Therefore, in this study, URLs, email addresses, and phone numbers were anonymized to avoid revealing such information.
Acknowledgements
We would like to thank Momose Oyama for the discussion and anonymous reviewers for their helpful advice. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110.
Code availability
Our code is available at https://github.com/ymgw55/Axis-Tour.
Appendix A Details of scatterplots from two-dimensional projection
This section explains the two-dimensional projection method used for the scatterplots in Fig. 1. We then present similar scatterplots for the three examples in Table 1. Finally, we define the metrics used in Section 1 to evaluate the quality of the scatterplots.
A.1 Scatterplot drawing method
We will explain the scatterplot drawing method using the Axis Tour embeddings.
First, we define a set of axis indices for projection. Let $I:=\{a,\ldots,b\}\,(a,b\in[d],a<b)$ be a consecutive interval of indices. The number of indices in $I$ is $|I|=b-a+1$. For example, in Fig. 1, $a=86$ and $b=94$.
Next, we define a matrix by extracting the axes of $\hat{\mathbf{T}}$ corresponding to $I$, where $\hat{\mathbf{T}}$ is the matrix obtained by normalizing the matrix $\mathbf{T}$. We denote this extracted matrix as $\hat{\mathbf{T}}_{I}\in\mathbb{R}^{n\times|I|}$.
[…]
$\mathbf{P}_{I}:=\begin{bmatrix}\varphi_{a}^{\top}\\ \vdots\\ \varphi_{b}^{\top}\end{bmatrix}\in\mathbb{R}^{|I|\times 2},$ (6)
where
$\varphi_{\ell}:=(\cos{\theta_{\ell}},\sin{\theta_{\ell}})^{\top}\in\mathbb{R}^{2},$ (7)
where
$\theta_{\ell}:=\frac{(\ell-a)\pi}{b-a}.$ (8)
Then we get the two-dimensional projection as $\mathbf{Q}_{I}:=\hat{\mathbf{T}}_{I}\mathbf{P}_{I}\in\mathbb{R}^{n\times 2}$. In $\mathbf{Q}_{I}$, the $\ell$-th axis of $\hat{\mathbf{T}}$ is projected along the direction of $\varphi_{\ell}$.
We denote the $i$-th word embedding of $\mathbf{Q}_{I}$ by $\mathbf{q}_{i}=(q_{i}^{x},q_{i}^{y})^{\top}\in\mathbb{R}^{2}$. When we plot the scatterplot of $\mathbf{Q}_{I}$, we do not plot the $i$-th word embedding if $q_{i}^{y}<0$ for visual clarity, and show the top five words for each axis. The indices of these top words equal to the following index set $\text{Show}_{I}$ defined with $\text{Top}_{k}^{\ell}$ from section 4.1:
$\text{Show}_{I}:=\{i\in[n]\mid q_{i}^{y}\geq 0\}\cap\bigcup_{\ell\in I}\text{Top}_{5}^{\ell}.$ (9)
Similarly, we can apply the same procedure to the Skewness Sort embeddings and obtain the two-dimensional scatterplot.
A.2 Scatterplots of Table 1
(c) The 23rd axis to the 31st axis
(d) The 101st axis to the 109th axis
(e) The 237th axis to the 245th axis
Figure 6: The scatterplots of the two-dimensional projections for the axes of the Axis Tour embeddings in Table 1. Similar to Fig. 1, we used the procedure described in Appendix A.1.
Figure 6 shows the scatterplots of the two-dimensional projections for the axes of the Axis Tour embeddings in Table 1, using the procedure described in Appendix A.1. Similar to Fig. 1, it is evident that the top words of the axes of the Axis Tour embeddings are farther from the origin than those of the Skewness Sort, and the meanings of the adjacent axes change continuously.
A.3 Evaluation metrics for scatterplots
In Section 1, we compared the quality of the scatterplots for Axis Tour and Skewness Sort by calculating the average distance of the top words from the origin. In this section, we first explain this metric and then, based on (4), define a new metric derived from the average of the cosine similarities between adjacent axis embeddings. We then compare these metrics for the scatterplots in Figs. 1 and 6. Similar to Appendix A.1, we use the Axis Tour embeddings to explain these metrics.
For the two-dimensional projection $\mathbf{Q}_{I}$ of the Axis Tour embeddings, we define the average distance $d_{I}$ of the top words from the origin, using the index set $\text{Show}_{I}$ in (9), as follows:
$d_{I}:=\frac{1}{|\text{Show}_{I}|}\sum_{i\in\text{Show}_{I}}\|\mathbf{q}_{i}\|$ (10)
A larger $d_{I}$ indicates that the scatterplot more accurately reflects the spatial distribution of the original embeddings, since the top words are positioned far from the origin.
We also define the average cosine similarity between adjacent axis embeddings, $c_{I}$, for the interval $I$ as follows:
$c_{I}:=\frac{1}{b-a}\sum_{\ell=a}^{b-1}\cos{(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})}$ (11)
As we saw in (4), since Axis Tour optimizes the order of the axes to maximize the sum of the cosine similarities, the value of $c_{I}$ reflects the semantic continuity of the axes in the scatterplot. It is important to note that $I$ represents a subinterval of $d$ indices, so we do not include the term of $\cos{(\mathbf{v}_{a},\mathbf{v}_{b})}$ in (11).
Both $d_{I}$ and $c_{I}$ can also be calculated in a similar manner for Skewness Sort. Table 2 shows the values of $d_{I}$ and $c_{I}$ for Figs. 1 and 6. In these examples, Axis Tour shows higher values for both $d_{I}$ and $c_{I}$ compared to Skewness Sort, indicating better projection quality.
In addition to the results for several subintervals in Table 2, we also see the semantic continuity of all axes. If $I=[d]$ in (11), we get $c_{I}=c_{[d]}$ as follows
$c_{[d]}=\frac{1}{d}\left\{\cos{(\mathbf{v}_{1},\mathbf{v}_{d})}+\sum_{\ell=1}^{d-1}\cos{(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})}\right\}$ (12)
where the term of $\cos{(\mathbf{v}_{1},\mathbf{v}_{d})}$ is added according to (4). The value of $c_{[d]}$ is equal to the average of the histogram in Fig. 2, so the value for Axis Tour is $0.244$ and for Skewness Sort is $0.017$, thus confirming the semantic continuity across all axes.
Axis Tour Skewness Sort
Fig. $d_{I}$ $c_{I}$ $d_{I}$ $c_{I}$
1 0.76 0.27 0.61 0.18
6 0.69 0.28 0.61 0.10
6 0.80 0.30 0.62 0.15
6 0.86 0.38 0.76 0.23
Table 2: The values of $d_{I}$ and $c_{I}$ for Figs. 1 and 6.
Appendix B Dimensionality reduction
This section details and supplements Section 4.3.
B.1 Definition of $\mathbf{f}_{r}$
Here we first discuss the definition of $\mathbf{f}_{r}$ in terms of skewness and then explain the normalization.
B.1.1 Skewness as weight
In (3), a vector $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for dimensionality reduction is defined by the skewness $\gamma_{\ell}\in\mathbb{R}_{\geq 0}$ of the $\ell$-th axis of $\mathbf{T}$. In particular, $f_{r}^{(\ell)}$, which corresponds to the weight of the $\ell$-th axis in the projection, is proportional to $\gamma_{\ell}^{\alpha}\,(\alpha\in\mathbb{R}_{\geq 0})$.
This is based on the assumption that the axis becomes more meaningful as the skewness increases [^13] While not specific to ICA-transformed embeddings, it is known from a study of sparse coding for language models that skewness correlates with interpretability Cunningham et al. (2023).. Higher-order statistics such as skewness are known to be sensitive to outliers Hyvärinen et al. (2001b), so we mitigate this effect by raising $\gamma_{\ell}$ to the power of $\alpha$. Since skewness is a third-order moment, we treat $\alpha=1/3$ as the default value. $\alpha=0$ is a uniform weight, while $\alpha=1$ is the original $\gamma_{\ell}$. Next, we will explain why $\mathbf{f}_{r}$ is normalized.
B.1.2 Reason for Normalizing $\mathbf{f}_{r}$
$\mathbf{f}_{r}$ is normalized (i.e., $\|\mathbf{f}_{r}\|=1$) to equalize the scale of the one-dimensional projections for each subspace. Thus, when $p=d$, $\mathbf{T}\mathbf{F}=\mathbf{T}$ holds.
For $p=d$, $I_{r}=\{r\}$, and from (3), $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}$ satisfies:
$f_{r}^{(\ell)}=\delta_{\ell r}=\begin{cases}1&\ell=r\\ 0&\ell\neq r\end{cases}.$ (13)
Here, $\delta_{\ell r}$ is the Dirac delta function. From (13) the matrix $\mathbf{F}=[\mathbf{f}_{1},\ldots,\mathbf{f}_{d}]\in\mathbb{R}^{d\times d}$ corresponds to the $d$-dimensional identity matrix $\mathbf{I}\in\mathbb{R}^{d\times d}$, so $\mathbf{T}\mathbf{F}=\mathbf{T}\mathbf{I}=\mathbf{T}$.
B.2 Projection from subspace to one-dimensional space
Figure 7: Projection of the subspace spanned by three consecutive axes in Fig. 1 into a one-dimensional space. Each word is assigned the color of the axis with the highest value. The projection direction is in the direction representing the subspace. For visualization, we randomly sampled 10,000 words, excluding the top five words on each axis.
This section explains the projection from a subspace to a one-dimensional space using a specific example. Consider the subspace spanned by three consecutive axes (89, 90, 91) from Fig. 1. Figure 7 shows the projection of this subspace using $\mathbf{f}_{r}$. The projection direction is in the direction representing the subspace, and the top words of each axis are projected close together.
Appendix C Distribution of cosine similarity
Let us consider two random vectors $X=(X_{1},\ldots,X_{d}),Y=(Y_{1},\ldots,Y_{d})\in\mathbb{R}^{d}$ with elements of mean zero $\mathbb{E}(X_{\ell})=\mathbb{E}(Y_{\ell})=0$ and finite variance $\mathbb{E}(X_{\ell}^{2})=\sigma_{X}^{2}$, $\mathbb{E}(Y_{\ell}^{2})=\sigma_{Y}^{2}$. We assume that the elements $X_{1},\ldots,X_{d}$ and $Y_{1},\ldots,Y_{d}$ are independent, and the sequence $X_{1}Y_{1},\ldots,X_{d}Y_{d}$ satisfies Lindeberg’s condition Billingsley (1995); for $Z_{\ell}=X_{\ell}Y_{\ell}/\sigma_{X}\sigma_{Y}$, $\forall\epsilon>0,\lim_{d\to\infty}d^{-1}\sum_{\ell=1}^{d}\mathbb{E}\Bigl{[}Z_{\ell}^{2}1(|Z_{\ell}|>\epsilon\sqrt{d})\Bigr{]}=0$. Lindeberg’s condition means no one $X_{\ell}Y_{\ell}$ dominates the inner product $\langle X,Y\rangle=\sum_{\ell=1}^{d}X_{\ell}Y_{\ell}$, and it is satisfied, for example, when all the elements follow the identical distribution.
Then, for sufficiently large $d$, the cosine similarity $\cos(X,Y)$ asymptotically follows $\mathcal{N}(0,1/d)$, the normal distribution with mean 0 and variance $1/d$. In other words, for sufficiently large $d$,
$\sqrt{d}\cos(X,Y)\sim\mathcal{N}(0,1).$ (14)
This is easily shown as follows. First note that $\mathbb{E}(X_{\ell}Y_{\ell})=\mathbb{E}(X_{\ell})\mathbb{E}(Y_{\ell})=0$, $\mathbb{E}(X_{\ell}^{2}Y_{\ell}^{2})=\mathbb{E}(X_{\ell}^{2})\mathbb{E}(Y_{\ell}^{2})=\sigma_{X}^{2}\sigma_{Y}^{2}$. Thus the inner product, if scaled by dimension, $d^{-1/2}\langle X,Y\rangle=d^{-1/2}\sum_{\ell=1}^{d}X_{\ell}Y_{\ell}$ has mean zero and variance $\sigma_{X}^{2}\sigma_{Y}^{2}$. Furthermore, according to the Lindeberg-Feller Central Limit Theorem, the distribution of the inner product asymptotically converges to the normal distribution as $d$ grows large:
$d^{-1/2}\langle X,Y\rangle\sim\mathcal{N}(0,\sigma_{X}^{2}\sigma_{Y}^{2}).$ (15)
It also follows from the law of large numbers, $d^{-1}\|X\|^{2}=d^{-1}\sum_{\ell=1}^{d}X_{\ell}^{2}$ converges in probability to $\mathbb{E}(X_{\ell}^{2})=\sigma_{X}^{2}$. Similarly $d^{-1}\|Y\|^{2}\to\sigma_{Y}^{2}$ in probability. Therefore,
$\sqrt{d}\cos(X,Y)=\frac{d^{-1/2}\langle X,Y\rangle}{\sqrt{d^{-1}\|X\|^{2}}\sqrt{d^{-1}\|Y\|^{2}}}$
converges to $d^{-1/2}\langle X,Y\rangle/\sigma_{X}\sigma_{Y}$ in probability, and thus (15) gives (14).
Appendix D Details of dimensionality reduction experiments in Section 5
$p=5$ $p=20$ $p=100$ $p=300$ Tasks PCA Rand. Skew. Tour. PCA Rand. Skew. Tour. PCA Rand. Skew. Tour. All Analogy capital-common-countries 0.00 0.00 0.00 0.00 0.37 0.00 0.06 0.11 0.95 0.56 0.85 0.87 0.95 capital-world 0.00 0.00 0.00 0.00 0.26 0.02 0.05 0.11 0.90 0.73 0.80 0.82 0.95 city-in-state 0.00 0.00 0.00 0.00 0.03 0.00 0.02 0.04 0.42 0.28 0.24 0.40 0.67 currency 0.00 0.00 0.00 0.00 0.02 0.00 0.00 0.02 0.09 0.08 0.08 0.10 0.12 family 0.01 0.00 0.00 0.00 0.40 0.02 0.08 0.22 0.78 0.68 0.80 0.75 0.88 gram1-adjective-to-adverb 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.01 0.08 0.08 0.14 0.09 0.21 gram2-opposite 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.14 0.12 0.18 0.14 0.26 gram3-comparative 0.00 0.00 0.00 0.00 0.05 0.01 0.04 0.11 0.62 0.46 0.58 0.66 0.88 gram4-superlative 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.01 0.31 0.54 0.23 0.31 0.69 gram5-present-participle 0.00 0.00 0.00 0.00 0.03 0.00 0.03 0.07 0.44 0.30 0.59 0.58 0.69 gram6-nationality-adjective 0.00 0.00 0.00 0.00 0.57 0.07 0.22 0.43 0.91 0.88 0.91 0.88 0.93 gram7-past-tense 0.00 0.00 0.00 0.00 0.05 0.02 0.04 0.06 0.45 0.36 0.47 0.51 0.60 gram8-plural 0.00 0.00 0.00 0.00 0.11 0.01 0.04 0.07 0.73 0.40 0.59 0.56 0.76 gram9-plural-verbs 0.00 0.00 0.00 0.00 0.06 0.01 0.02 0.06 0.39 0.27 0.29 0.53 0.58 jj_jjr 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.03 0.35 0.23 0.29 0.43 0.66 jj_jjs 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.01 0.21 0.36 0.14 0.20 0.51 jjr_jj 0.00 0.00 0.00 0.00 0.01 0.00 0.03 0.02 0.33 0.27 0.32 0.33 0.54 jjr_jjs 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.01 0.24 0.37 0.15 0.20 0.55 jjs_jj 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.01 0.21 0.13 0.18 0.19 0.48 jjs_jjr 0.00 0.00 0.00 0.00 0.01 0.00 0.01 0.01 0.29 0.16 0.25 0.33 0.63 nn_nnpos 0.00 0.00 0.00 0.00 0.05 0.01 0.01 0.04 0.35 0.20 0.29 0.28 0.42 nn_nns 0.00 0.00 0.00 0.00 0.04 0.00 0.04 0.06 0.55 0.33 0.49 0.51 0.74 nnpos_nn 0.00 0.00 0.00 0.00 0.03 0.00 0.02 0.06 0.40 0.20 0.34 0.31 0.45 nns_nn 0.00 0.00 0.00 0.00 0.05 0.00 0.04 0.06 0.48 0.30 0.43 0.44 0.64 vb_vbd 0.00 0.00 0.00 0.00 0.11 0.03 0.09 0.07 0.45 0.36 0.40 0.54 0.58 vb_vbz 0.00 0.00 0.00 0.00 0.09 0.02 0.05 0.08 0.58 0.33 0.50 0.68 0.76 vbd_vb 0.00 0.00 0.00 0.00 0.09 0.04 0.12 0.08 0.46 0.36 0.56 0.56 0.69 vbd_vbz 0.00 0.00 0.00 0.00 0.08 0.01 0.04 0.05 0.48 0.26 0.38 0.54 0.63 vbz_vb 0.00 0.00 0.00 0.00 0.06 0.02 0.06 0.11 0.65 0.43 0.70 0.70 0.82 vbz_vbd 0.00 0.00 0.00 0.00 0.05 0.01 0.03 0.05 0.33 0.33 0.42 0.44 0.55 Average 0.00 0.00 0.00 0.00 0.09 0.01 0.04 0.07 0.45 0.34 0.42 0.46 0.63 Similarity MEN 0.16 0.19 0.11 0.35 0.32 0.33 0.29 0.51 0.66 0.56 0.63 0.66 0.75 MTurk 0.17 0.08 0.12 0.32 0.38 0.32 0.30 0.52 0.57 0.53 0.57 0.61 0.64 RG65 0.31 0.09 0.05 0.29 0.36 0.42 0.28 0.50 0.68 0.66 0.59 0.63 0.78 RW 0.09 0.10 0.07 0.13 0.14 0.16 0.10 0.25 0.24 0.32 0.28 0.30 0.34 SimLex999 0.01 0.13 0.04 0.07 0.11 0.21 0.08 0.21 0.27 0.37 0.28 0.31 0.40 WS353 0.12 0.16 0.02 0.31 0.15 0.28 0.18 0.44 0.47 0.40 0.43 0.52 0.57 WS353R 0.12 0.14 0.01 0.17 0.15 0.16 0.15 0.35 0.40 0.28 0.35 0.44 0.51 WS353S 0.18 0.21 0.06 0.45 0.21 0.35 0.26 0.55 0.57 0.51 0.58 0.62 0.69 Average 0.15 0.14 0.06 0.26 0.23 0.28 0.20 0.42 0.48 0.45 0.46 0.51 0.59 Categorization AP 0.33 0.22 0.22 0.27 0.36 0.26 0.28 0.40 0.51 0.45 0.54 0.56 0.66 BLESS 0.31 0.28 0.27 0.36 0.42 0.36 0.35 0.51 0.73 0.68 0.69 0.76 0.79 Battig 0.18 0.10 0.12 0.15 0.24 0.14 0.16 0.22 0.37 0.29 0.35 0.34 0.42 ESSLI_1a 0.50 0.41 0.45 0.64 0.61 0.57 0.48 0.77 0.73 0.59 0.73 0.75 0.70 ESSLI_2b 0.47 0.62 0.45 0.53 0.73 0.68 0.55 0.68 0.75 0.75 0.70 0.75 0.78 ESSLI_2c 0.44 0.44 0.42 0.44 0.53 0.60 0.49 0.56 0.58 0.53 0.60 0.60 0.58 Average 0.37 0.35 0.32 0.40 0.48 0.44 0.38 0.52 0.61 0.55 0.60 0.63 0.65
Table 3: The performance of dimensionality reduction for $p$-dimensional embeddings. Rand. stands for Random Order, Skew. for Skewness Sort, and Tour. for Axis Tour. The values in the table correspond to top 1 accuracy for analogy tasks, Spearman’s rank correlation for word similarity tasks, and purity for categorization tasks. Note that at $p=300$, all embeddings give the same results.
Figure 8: The performance of dimensionality reduction for the PCA-transformed embeddings and the Axis Tour embeddings with $\alpha=0,1/3,1/2,1$. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks.
D.1 Detailed explanation of each task
Analogy task.
The embedding of the $\text{word}_{i}$ is denoted by $\mathbf{y}_{i}\in\mathbb{R}^{d}$. We used the Google Analogy Test Set Mikolov et al. (2013a), which contains 14 types of analogy tasks, and the Microsoft Research Syntactic Analogies Dataset Mikolov et al. (2013c), which contains 16 types of analogy tasks. In the analogy tasks, the quality of the embeddings is evaluated by inferring $\text{word}_{4}$ to which $\text{word}_{3}$ corresponds if $\text{word}_{1}$ corresponds to $\text{word}_{2}$. We compute the vector $\mathbf{y}_{2}-\mathbf{y}_{1}+\mathbf{y}_{3}$ and see if the closest embedding is $\mathbf{y}_{4}$ (top1 accuracy).
Word similarity task.
We used MEN Bruni et al. (2014), MTurk Radinsky et al. (2011), RG65 Rubenstein and Goodenough (1965), RW Luong et al. (2013), SimLex999 Hill et al. (2015), WS353 Finkelstein et al. (2002), WS353R (WS353 Relatedness), and WS353S (WS353 Similarity). In the word similarity tasks, the quality of the embeddings is evaluated by measuring the cosine similarity of the word embeddings and comparing it to the human-rated similarity scores. As the evaluation metric, we used Spearman’s rank correlation coefficient between the human ratings and the cosine similarity.
Categorization task.
We used AP Almuhareb and Poesio (2005), BLESS Baroni and Lenci (2011), Battig Battig and Montague (1969), ESSLLI_1a Baroni et al. (2008), ESSLLI_2b Baroni et al. (2008), and ESSLLI_2c Baroni et al. (2008). In the categorization tasks, the quality of the embeddings is evaluated by clustering them in the setting where each word is assigned a class label. As the evaluation metric, we used Purity, which shows the proportion of the most frequent class in the clusters. As clustering methods, we used Hierarchical Clustering with five settings [^14] By default, Word Embedding Benchmark uses the following affinity and linkage pairs for hierarchical clustering: (affinity, linkage) = (euclidean, ward), (euclidean, average), (euclidean, complete), (cosine, average), (cosine, complete). and $K$-means [^15] We used the same seed for all experiments., and then selected the one that gave the highest purity.
D.2 Results
Table 3 shows detailed experimental results of PCA, Random Order, Skewness Sort, and Axis Tour at $p=5,20,100,300$ for each task. As already seen, Fig. 4 in Section 5 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the embeddings.
The Axis Tour embeddings showed superior performance in the word similarity tasks and the categorization tasks for almost all dimensions compared to other methods. In the analogy tasks, the Axis Tour embeddings achieved performance comparable to PCA and better than Random Order and Skewness Sort in most dimensions.
D.3 Setting of $p=d\,(=300)$
Note that the experimental results are the same for all embeddings for $p=300$. First, as we saw in Appendix B.1, when $p=d\,(=300$), the matrix $\mathbf{TF}$ (i.e., the projected $p$-dimensional embeddings) is equal to the matrix $\mathbf{T}$ (i.e., the $d$-dimensional Axis Tour embeddings). Then, by definition, Axis Tour, Random Order, and Skewness Sort are the embeddings obtained by reordering the axes of the ICA-transformed embeddings and flipping their signs as needed. Thus, these three can be seen as the embeddings obtained by applying an orthogonal matrix to the ICA-transformed embeddings. Since the ICA-transformed embeddings are derived from the PCA-transformed embeddings by applying an orthogonal matrix [^16] Refer to the previous work for the relationship between PCA and ICA Yamagiwa et al. (2023)., cosine similarity and Euclidean distance remain unchanged for PCA, Random Order, Skewness Sort, and Axis Tour, leading to identical results in downstream tasks.
Appendix E Additional experiments
E.1 Qualitative observations for all axes of the Axis Tour embeddings
0 1 2 3 4 5 6 7 8 9
phaen region mountain stage italy juan da state india vaas
sandretto goriška mount vinokourov italian spain são terengganu indian jayasuriya
nakhchivan languedoc mountains vuelta di luis paulo kedah singh wicket
burghardt regions everest stages francesco gonzález joão perlis shri jayawardene
regno saguenay-lac-saint-jean peaks magicians pietro garcía janeiro kelantan delhi kallis
10 11 12 13 14 15 16 17 18 19
andrew chief general bhavsar spearritt contributed by ***.com micro-history family
divoff executive then-attorney beppe 297.00 sebti sivuyile tburr 1977-2010 lythraceae
vanwyngarden ceo gen. ayme moçambique avet mulyanda fluto 1951-1972 chrysobalanaceae
hruska justiceship sindiso kahrd epoca aygin boonradom ***.com qc8 polyporaceae
hampsten hienonen jiyane ripstein huajun toosi tolkun ***.com 1977-2006 pyronemataceae
20 21 22 23 24 25 26 27 28 29
order the hungarians serb russian czech germany france canada australia
neuroptera macdougalls ethnic bosnian russia prague german french canadian australian
boletales powhatans asians croatia moscow poland berlin le ontario queensland
svu andhras tatars croatian sergei polish von paris quebec brisbane
poverelle sasanians berbers serbian aleksandr warsaw cologne du saskatchewan perth
30 31 32 33 34 35 36 37 38 39
wiltshire liga coach frank tony mike michael joseph jack &amp;
shrewsbury relegated coaching capra ianno johanns finnissy joe palance llp
lincolnshire fc scolari sinatra canzoneri fetters cerveris macenka o’lantern amp
peterborough f.c. vogts wisner oursler petke stuhlbarg papp spring-heeled &
croydon serie capello aigbogun kornheiser lupica tomasky nollekens lemmon firm
40 41 42 43 44 45 46 47 48 49
analyst an three-week festivities 27th striker goalkeeper mickelson 6-4 2-9
strategist rohp 20-minute celebration 22nd equalized keeper furyk 7-5 8-8
securities average-sized two-week celebrations 26th midfielder goalie els roddick 1-9
hyoty waidhofen two-month commemoration 23rd header goaltender faldo kafelnikov 4-12
udomsirikul taisce 15-minute ceremonies 35th equalised jaaskelainen woods dementieva 17-65
50 51 52 53 54 55 56 57 58 59
24-14 touchdown rebounds inning .301 9-for-13 on island greek george
24-17 quarterback hardaway hitter .292 3-for-8 premised islands greece takei
14-10 touchdowns pippen pitcher .293 5-for-12 picturized fuerteventura athens stroumboulopoulos
20-10 qb shaquille outfielder .289 4-for-11 ixnay conanicut greeks w.bush
27-17 interceptions mcdyess baseman .288 12-of-19 gorging wangerooge zeus maharis
60 61 62 63 64 65 66 67 68 69
president burundi envoys between brouhaha challenges numerous mistake horrific vitriolic
teburoro uganda talks relationship over dilemmas various error terrible racist
vice tanzania envoy quarrel editorship confronting sundry mistakes horrendous denouncing
bagbandy kenya annan rivalry dispute dilemma other misjudgments horrible strident
issayas zambia solana relationships damocles vexing mishaps disregard unspeakable insults
70 71 72 73 74 75 76 77 78 79
disgust tenacity informs life their arguably livelier conclusive endear slowing
anger humility destroys expectancies our liveliest nastier corroboration semblance slowdown
feelings toughness sustains commuted forbears costliest rougher no wobbled weakening
sadness newfound confronts transience fellow-citizens quietest prettier substantiation shaky sluggish
revulsion professionalism goes great-west its best-preserved deadlier scant unscathed decline
80 81 82 83 84 85 86 87 88 89
index 1.48 20.45 30-year 2 sauce beer antidepressant proteins chlorine
asx 1.62 15.55 yield 1 cooked drink drugs protein solvents
ase 2.07 33.65 bond teaspoons cheese drinks drug genes ammonia
ftse 2.05 13.45 10-year 3 roasted brewed medications gene liquid
klse 1.67 17.35 yields teaspoon bread drinkers prozac rna flammable
90 91 92 93 94 95 96 97 98 99
copper oil power line road three-story hotel stores industry subsidiary
nickel petroleum electricity railway highway brick hotels store export company
manganese oilfields substation lines route facade resort grocery manufacturing maker
molybdenum crude hydroelectric arbatsko-pokrovskaya north-south two-story resorts supermarkets distributors alcatel
ore pemex megawatts kolsås two-lane building marriott retailer shoemaking corp.
100 101 102 103 104 105 106 107 108 109
hedge pay land laws court lawsuits charges camp corpses remain
fund fees property regulations judge lawsuit alleged prison corpse remained
funds payments lands enacted appellate litigation prosecutors buchenwald exhumed stayed
investments payment estate law appeals suits indicted camps dismembered stubbornly
investing paid bergisches provisions supreme suit convicted inmates bodies stays
110 111 112 113 114 115 116 117 118 119
hopelessly incredibly ingenious to fostering tricking into has 19,583 3,048
frustrated amazingly devious intend initiatives busily morphed had 21,563 dolne
woefully extremely clever able sustainable concentrating transmuted beeen 16,875 prateek
hamstrung very intricate humiliate empowering classifying degenerates consistantly 20,833 bugis
chronically wonderfully cunning try entrepreneurship disposing delving reinvented 30,313 gumbinnen
Table 4: The top five words for the 00th axis to the $119$th axis when we apply Axis Tour to 300-dimensional GloVe.
120 121 122 123 124 125 126 127 128 129
system desktop phone web television 2300 4:23 _ new 121.58
systems macintosh cellular sites tv 12:30 2:47 ondeck york 114.78
renin-angiotensin pc cellphone site channel 11:30 2:33 holimont orleanians 121.22
centralized software telephone online broadcast 0330 8:38 hawksnest orleans-based 119.76
computerised server wireless myspace cbs 0930 2:21 belleayre n.y. 106.12
130 131 132 133 134 135 136 137 138 139
seventies 2001-2003 four 89 445 1,149 209.6 95.3 28.4 5.5
sixties 2003-2005 five 83 244 4,461 218.8 96.2 23.1 4.8
2010s 1999-2001 three 86 285 1,737 218.3 89.8 23.9 5.7
1800s 1998-2001 eleven 85 292 1,701 308.9 89.9 26.1 4.4
early 1995-1997 six 88 344 1,109 246.3 93.1 27.6 4.6
140 141 142 143 144 145 146 147 148 149
23,000 union alike theories church sex actress german-born old multibillion
110,000 workers academics theory anglican sexual wife newspaperman 35-year 20-million
39,000 machinists industrialists posited episcopal homosexual mother entrepreneur 14-year multi-million
43,000 unions sociologists notions congregations heterosexual daughter politician 50-year 100-million
48,000 teamsters pundits notion congregation unmarried née russian-born 22-year multimillion
150 151 152 153 154 155 156 157 158 159
) rights group force insurgents guns wheels sedan schumacher preakness
( human groups gendarmerie militants rifles wheel sedans barrichello belmont
.0358 aprodeh lashkar-e-jhangvi contingent gunmen gun rudder v8 massa filly
3.7996 zimrights forzani 500-strong guerrillas caliber brakes turbocharged ferrari baffert
unitals pillay harkat contingents forces weapon hydraulic camry raikkonen thoroughbred
160 161 162 163 164 165 166 167 168 169
400m sports welterweight la que basidiomycota .2667 lb3 drawno behshahr
100m sport heavyweight cerva pero gorlice estrategia 5lb krośniewice abyek
200m softball middleweight cenerentola sus empleos creado 8lb brochów ramian
freestyle volleyball ibf ferrière en autovía hirta lb7 lubaczów khamir
200-meter soccer wba louvière una pudiera a.m.-6 4lb pobiedziska javanrud
170 171 172 173 174 175 176 177 178 179
bank river lake city international thailand minister spokesman nafez adtac
central ljubljanica mývatn ozamis airport thai prime faizasyah sedwill coronae
nivard tigris erhai malaybalay tocumen nakhon minster cirtek british-controlled maḩalleh
banks tributary chilka phenix gökçen bangkok minsiter tsanchev videoton decarboxylase
pboc rivers waramaug hitec aiport thais interior ladsous 2-6-2 poderoso
180 181 182 183 184 185 186 187 188 189
8a-4p qeshlaqi automoviles jreyes ***-***-**** _____… moodin - :
romik dizaj-e vx-6 ilovar .0210 interbk ***-***-**** skway bkh www.***.com
shefer kalayeh ***-***-**** hohtz ***-***-**** ooooo… wehz thoh wc2006-asia http://www.***.org
samayoa now-e principalmente ray-finned ***@***.com ***-***-**** prahnk kursh manutd http://www.***.com
hermanas patryan yengejeh odalovic 65stk harrynytimes eesah tsih pickup5thgraf http://www.***.org
190 191 192 193 194 195 196 197 198 199
; sarā rd2 ipnotinmx gibbosa 1507.50 ixmiquilpan prolegs masuku artayev
priu as clientes mordella pratylenchus .000663 ***@***.com grij quartic autuori
chahār weinzapfel significado bifrenaria analista .71078 macul havner tayyab
.8226 khvosh desempleo sandwicense fimbriatus ramnarine jagdeep rambusch shealy pareto
sa/b akbarābād gevar ***-***-**** laticeps hàm naab yr.ago vidro zermelo
200 201 202 203 204 205 206 207 208 209
swaffham phthalic 72-77 magnifica funderburke cunxiao wang japan south park
**-***-**** chamba gildernew compra hintikka cunshen zhang japanese korean prenton
ollivier nguyên deanne qc7 chofetz cunxu xu tokyo korea ji-sung
zend kinghorn ethelbert kowalska danys siyuan li akira dakotans naturpark
cronulla-sutherland matanuska-susitna kether handanovic opticians lanqing liu takashi carolinians breffni
210 211 212 213 214 215 216 217 218 219
national parliament party polling republican . i silently trousers whitish
winema duma janata electoral sen. , ’cause stared pants yellowish
jurnalul speaker bjp balloting gop but gonna screamed dresses brownish
ranthambhore 150-seat socialists election republicans though yeah yelled dress greyish
chúa 500-seat ndp tallying mccain even gosh sobbing wearing reddish-brown
220 221 222 223 224 225 226 227 228 229
ultraviolet spacecraft aircraft vessel hurricane temperatures grasslands shrubs mammals virus
infrared astronauts jet ship storm humid habitats trees birds h5n1
telescopes astronaut planes ships storms chilly soils vines rabbits swine
light nasa boeing vessels typhoon weather marshes planted animals flu
wavelengths orbit f-15 boats cyclone unseasonably sediments seedlings reptiles outbreak
230 231 232 233 234 235 236 237 238 239
diagnosed medical undergo survey agency newspaper ign award film superhero
lung care thorough surveys notimex daily popmatters awards films marvel
inflammation hospital undergoing statistics kathpress zeitung allmusic awarded movie spin-off
complications hospitals evaluation gallup bss izvestia reviewer prize starring superheroes
fractures physician undergone statistical telam kommersant gamespot emmy directed characters
Table 5: The top five words for the $120$th axis to the $239$th axis when we apply Axis Tour to 300-dimensional GloVe.
240 241 242 243 244 245 246 247 248 249
album piano paintings manuscript language name formula_1 set mark brian
albums violin painting biographies languages names formula_2 aflame high-water trenchard-smith
band cello art pages pashto surname formula_3 10-cd dindal doyle-murray
self-titled percussion sculpture book colloquial phrase formula_4 setting wainberg cadd
ep orchestral watercolor handwritten dialect misspelling formula_5 4-cd hoppus donlevy
250 251 252 253 254 255 256 257 258 259
ireland martin peter paul john bazzani david scott 2-54
belfast scelzo guillam virilio rhys-davies looking-glass pittu livengood cretier
irish wansleben shockheaded ricoeur canemaker xff mouret proval lobdell veltman
mowlam clunes maffay langmack motson munro margolick gud f.r.
northern damm rauhofer celan mcgahern schizopolis mármol gorcyca speedman nahant
260 261 262 263 264 265 266 267 268 269
palm county calif. school university 1976 1853 february king alwaleed
beach unincorporated california high professor 1973 1852 october sigismund prince
fla. dekalb inglewood elementary graduate 1966 1847 december iv saudi
ostrowski fayette pasadena pine-richland doctorate 1968 1856 june emperor kingdom
broward pulaski pomona jr/sr faculty 1971 1854 april iii tupouto’a
270 271 272 273 274 275 276 277 278 279
ali israel daniel swedish thomas william charles james wilkison nack
al israeli briere sweden kretschmann mastrosimone dutoit frain mudavadi narrowly
mohammed israelis macivor norwegian aquinas beaudine grodin roday henwood gyoergy
sheikh netanyahu gildenlöw norway quasthoff fichtner sheeler luceno hogzilla khand
mohammad aviv balavoine fredrik fingar saroyan wuorinen remar mottola vanak
280 281 282 283 284 285 286 287 288 289
nhls janyk vranjes robert van chris marchena steve shaara richard
ba’asyir tookie kilvert deniro lieshout volstad staniforth railsback jeff stoltzman
organizaciones germanica bukan halmi dijk braide kaboul turre shain ayoade
yordanov walthers paratore garrigus zandt bordano pelzer forbert friesen basehart
millán adquisiciones ponge redford tuyl tidland 43.18 kimock sharlet sandomir
290 291 292 293 294 295 296 297 298 299
simon malatesta bilyaletdinov .000106 nazione olya gearon sirajul wonk kerberos
boccanegra .0217 pouget fhs ba872 dodecahedral mikuláš overexcited kappa mckelvey
napier-bell lockard cytidine colecovision blouin cih bracigliano nabhani fraternity wajir
vouet 330-pound zhulali ronghua comunicacion alnus venero then-reigning godsmack veg
callow puji bc4 yunlong wachau 9.29 mariangela karmichael bibliophiles pdca
Table 6: The top five words for the $240$th axis to the $299$th axis when we apply Axis Tour to 300-dimensional GloVe.
In Table 1, we used 300-dimensional GloVe and showed the semantic continuity of the axes of the Axis Tour embeddings, with illustrative examples of the meaning of countries (the $23$rd axis to the $31$st axis), law (the $101$st axis to the $109$th axis), and art (the $237$th axis to the $245$th axis). This section presents the top five words of the normalized embeddings across all 300 axes in Tables 45,and 6.
For example, in Table 4, the $45$th and $46$th axes are related to soccer, the $47$th axis to golf, the $48$th axis to tennis, the $49$th and $50$th axes to scores, the $51$st axis to American football, the $52$nd axis to basketball, and the $53$rd axis to baseball. These axes illustrate the semantic continuity across sports.
In Table 5, the $129$th to the $140$th axes are related to numbers. It is also interesting to note that the top words of each axis have similar numerical scales, highlighting how well the axis of the ICA-transformed embeddings captures meaning. In addition, the meaning of each axis changes continuously, much like a game of word association: the $219$th axis relates to colors, the 220th axis to light, the 221st axis to space, the 222nd axis to airplanes, the 223rd axis to ships, the 224th axis to storms, the 225th axis to weather, the 226th axis to biomes, the 227th axis to plants.
In Table 6, the 272nd to the 289th axes are related to personal names from different linguistic regions as if this were a cluster of the meaning.
Note that due to space limitations, the top 1 and top 3 words on the $185$th axis are truncated because they are repetitive symbols, and that URLs, email addresses, and phone numbers are anonymized.
E.2 Comparison of $\alpha$
Figure 8 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the PCA-transformed embeddings and the Axis Tour embeddings with $\alpha=0,1/3,1/2,1$.
From Fig. 8, we can see that the performance of the Axis Tour embeddings changes for each task, depending on $\alpha$. For example, when comparing across all $\alpha$, while $\alpha=0$ shows good performance on word similarity tasks and poor performance on analogy tasks, $\alpha=1$ shows the opposite.
These results suggest that the quality of low-dimensional embeddings by the Axis Tour embeddings depends on the vector $\mathbf{f}_{r}$ for projection. However, the overall changes for each $\alpha$ are not as large, and in all tasks the performance is better than or comparable to that of the PCA-transformed embeddings, indicating the ability to construct better or comparable low-dimensional embeddings compared to PCA.
E.3 Comparisons of $k$
While we have done experiments for the Axis Tour with $k=100$ using 300-dimensional Glove in Section 5, when computing the axis embedding in (3), what is the appropriate value for $k$?
For example, if $k=1$ and the top $1$ word happens to have a meaning different from that of the axis, it is not desirable to define the axis embedding using only that word. Conversely, as $k$ increases, the number of words with meanings different from that of the axis in the top $k$ words also increases, hindering the ability of the axis embedding to represent the meaning of the axis. For example, in an extreme case where $k=n$, $\text{Top}_{k}^{\ell}$ becomes equal to $[n]$, and all axis embeddings become the mean vector of $\hat{\mathbf{s}}_{i}$ over the vocabulary set. In this case, it impossible to find the order that maximizes the semantic continuity of the axes.
In this section, to address these questions, we compare the metric for the semantic continuity of the axes, as defined in (4), for $k=1,10,100,1000$, and then perform qualitative observation and dimension reduction experiments.
E.3.1 Selection of $k$
In preparation, we perform the Axis Tour for $k_{1}=1,10,100,1000$, resulting in the embedding matrices $\mathbf{T}_{1},\mathbf{T}_{10},\mathbf{T}_{100},\mathbf{T}_{1000}$. In this section, we redefine the axis embedding for them with top $k_{2}=1,10,100,1000$, then evaluate the metric for the semantic continuity of the axes and thus compare the quality of $\mathbf{T}_{k_{1}}$.
Redefinition of axis embedding.
Similar to $\hat{\mathbf{S}}$, we define the matrix $\hat{\mathbf{T}}_{k_{1}}\in\mathbb{R}^{n\times d}$ as the normalization of the embeddings $\mathbf{T}_{k_{1}}$ with row vectors $\hat{\mathbf{t}}_{k_{1},i}=\mathbf{t}_{k_{1},i}/\|\mathbf{t}_{k_{1},i}\|$. Here, the $i$-th word embedding of $\mathbf{T}_{k_{1}}$ and $\hat{\mathbf{T}}_{k_{1}}$ are denoted by $\mathbf{t}_{k_{1},i},\hat{\mathbf{t}}_{k_{1},i}\in\mathbb{R}^{d}$, respectively. We compare the elements of the $\ell$-th axis of $\hat{\mathbf{T}}_{k_{1}}$ and denote the index set of words corresponding to the top $k_{2}$ elements as $\text{Top}_{k_{2}}^{\ell}$.
We redefine the $\ell$-th axis embedding $\mathbf{v}_{\ell}{(k_{1},k_{2})}$ for $k_{1}$ and $k_{2}$ as follows:
$\mathbf{v}_{\ell}{(k_{1},k_{2})}:=\frac{1}{k_{2}}\sum_{i\in\text{Top}_{k_{2}}^{\ell}}\hat{\mathbf{t}}_{k_{1},i}\in\mathbb{R}^{d}.$ (16)
If $k_{2}=k_{1}$, then $\mathbf{v}_{\ell}{(k_{1},k_{2})}$ coincides with the axis embedding used in the Axis Tour optimization. In contrast, if $k_{2}\neq k_{1}$ and the semantic continuity of the axes is still observed for $\mathbf{v}_{\ell}{(k_{1},k_{2})}$, then $\mathbf{T}_{k_{1}}$ can be considered as high-quality embeddings that are robust to changes for $k_{2}$.
Semantic continuity of axes for $\mathbf{v}_{\ell}{(k_{1},k_{2})}$.
Using $\mathbf{v}_{\ell}{(k_{1},k_{2})}$, we define the metric for the semantic continuity of the axes for each $k_{1}$ and $k_{2}$ as follows:
$c(k_{1},k_{2}):=\frac{1}{d}\cos(\mathbf{v}_{1}{(k_{1},k_{2})},\mathbf{v}_{d}{(k_{1},k_{2})})+\\ \frac{1}{d}\sum_{\ell=1}^{d-1}\cos(\mathbf{v}_{\ell}{(k_{1},k_{2})},\mathbf{v}_{\ell+1}{(k_{1},k_{2})}).$ (17)
Note that, unlike (4), the axes are already ordered by the Axis Tour using $k_{1}$. Furthermore, in (17), the sum of $\cos(\mathbf{v}_{\ell}{(k_{1},k_{2})},\mathbf{v}_{\ell+1}{(k_{1},k_{2})})$ is divided by the dimension $d$, which can be interpreted as the average of the cosine similarities between adjacent axis embeddings $\mathbf{v}_{\ell}{(k_{1},k_{2})}$ and $\mathbf{v}_{\ell+1}{(k_{1},k_{2})}$.
Comparison method for $k$.
We aim to compare the quality of $\mathbf{T}_{k_{1}}$. To facilitate this, we define the univariate functions of $k$ for $k_{1}=1,10,100,1000$:
$C_{1}(k)$ $:=c(1,k),$ (18)
$C_{10}(k)$ $:=c(10,k),$ (19)
$C_{100}(k)$ $:=c(100,k),$ (20)
$C_{1000}(k)$ $:=c(1000,k).$ (21)
For example, $C_{1}(k)$ in (18) is a function that measures the semantic continuity for the axis embedding $\mathbf{v}_{\ell}{(1,k)}$ redefined by the top $k$ words of the $\ell$-th axis of $\mathbf{T}_{1}$. To evaluate the robustness of $\mathbf{T}_{1}$ to changes for $k$, we average the values of $C_{1}(k)$ for $k=1,10,100,1000$. A higher average value indicates better quality.
The same explanation can be applied to $\mathbf{T}_{10}$, $\mathbf{T}_{100}$, $\mathbf{T}_{1000}$, which leads to the definition of a function $M(k)$ for calculating these averages:
$M(k):=\\ \frac{C_{k}(1)+C_{k}(10)+C_{k}(100)+C_{k}(1000)}{4}$ (22)
We compare $M(1),M(10),M(100),M(1000)$ to evaluate the quality of $\mathbf{T}_{1},\mathbf{T}_{10},\mathbf{T}_{100},\mathbf{T}_{1000}$.
Results.
Figure 9: Plots of the functions $C_{1}(k)$, $C_{10}(k)$, $C_{100}(k)$, $C_{1000}(k)$ and $M(k)$ for $k=1,10,100,1000$.
237 238 239 240 241 242 243 244 245
india thailand japan germany france italy ireland czech hungarians
indian thai japanese german french italian belfast prague ethnic
singh nakhon tokyo berlin le di irish poland asians
shri bangkok akira von paris francesco mowlam polish tatars
delhi thais takashi cologne du pietro northern warsaw berbers
112 113 114 115 116 117 118 119 120
to order court charges lawsuits challenges remain hopelessly incredibly
intend neuroptera judge alleged lawsuit dilemmas remained frustrated amazingly
able boletales appellate prosecutors litigation confronting stayed woefully extremely
humiliate svu appeals indicted suits dilemma stubbornly hamstrung very
try poverelle supreme convicted suit vexing stays chronically wonderfully
220 221 222 223 224 225 226 227 228
sedan van paintings manuscript piano album award actress film
sedans lieshout painting biographies violin albums awards wife films
v8 dijk art pages cello band awarded mother movie
turbocharged zandt sculpture book percussion self-titled prize daughter starring
camry tuyl watercolor handwritten orchestral ep emmy née directed
(a) $k=1$
102 103 104 105 106 107 108 109 110
6-4 mickelson schumacher stage france italy da juan welterweight
7-5 furyk barrichello vinokourov french italian são spain heavyweight
roddick els massa vuelta le di paulo luis middleweight
kafelnikov faldo ferrari stages paris francesco joão gonzález ibf
dementieva woods raikkonen magicians du pietro janeiro garcía wba
199 200 201 202 203 204 205 206 207
between sex laws court lawsuits charges camp corpses rights
relationship sexual regulations judge lawsuit alleged prison corpse human
quarrel homosexual enacted appellate litigation prosecutors buchenwald exhumed aprodeh
rivalry heterosexual law appeals suits indicted camps dismembered zimrights
relationships unmarried provisions supreme suit convicted inmates bodies pillay
112 113 114 115 116 117 118 119 120
preakness award ign album piano paintings manuscript web phone
belmont awards popmatters albums violin painting biographies sites cellular
filly awarded allmusic band cello art pages site cellphone
baffert prize reviewer self-titled percussion sculpture book online telephone
thoroughbred emmy gamespot ep orchestral watercolor handwritten myspace wireless
(b) $k=10$
104 105 106 107 108 109 110 111 112
the river lake canada france la italy stage piano
macdougalls ljubljanica mývatn canadian french cerva italian vinokourov violin
powhatans tigris erhai ontario le cenerentola di vuelta cello
andhras tributary chilka quebec paris ferrière francesco stages percussion
sasanians rivers waramaug saskatchewan du louvière pietro magicians orchestral
189 190 191 192 193 194 195 196 197
fostering tricking laws court lawsuits charges camp set formula_1
initiatives busily regulations judge lawsuit alleged prison aflame formula_2
sustainable concentrating enacted appellate litigation prosecutors buchenwald 10-cd formula_3
empowering classifying law appeals suits indicted camps setting formula_4
entrepreneurship disposing provisions supreme suit convicted inmates 4-cd formula_5
108 109 110 111 112 113 114 115 116
france la italy stage piano album superhero film
french cerva italian vinokourov violin albums marvel films
le cenerentola di vuelta cello band spin-off movie xff
paris ferrière francesco stages percussion self-titled superheroes starring
du louvière pietro magicians orchestral ep characters directed schizopolis
(c) $k=1000$
Table 7: Semantic continuity of axes by Axis Tour for normalized ICA-transformed embeddings. First, we focus on each central axis of Table 1 where $k=100$. The axes are the 27th axis (france, french, le, paris, du), the 105th axis (lawsuits, lawsuit, litigation, suits, suit), and the 241st axis (piano, violin, cello, percussion, orchestral). This table then shows the top five words of the axes near these selected axes for $k=1,10,1000$. Note that the axis indices change depending on the results of each Axis Tour.
Figure 10: The performance of dimensionality reduction for the PCA-transformed embeddings and the Axis Tour embeddings with $k=1,10,100,1000$ and $\alpha=1/3$. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks.
Figure 9 shows plots of the functions $C_{1}(k)$, $C_{10}(k)$, $C_{100}(k)$, $C_{1000}(k)$ and $M(k)$ for $k=1,10,100,1000$. $M(100)$ is the maximum value of $M(k)$. This result validates the use of $k=100$ as the default value for our experiment settings.
Furthermore, $C_{1}(1)$ is the maximum value among $C_{1}(k)$, $C_{10}(k)$, $C_{100}(k)$, $C_{1000}(k)$ $(k=1,10,100,1000)$. In this setting, since we use the top 1 word of each axis for axis embedding, this Axis Tour is equivalent to the Word Tour of 300 words using cosine similarity distance. Thus, the axis embedding is identical to the word embedding, which avoids ambiguity and simplifies the task of finding semantically similar words. However, the lower value of $M(1)$ compared to $M(10),M(100),M(1000)$ illustrates the instability of representing the meaning of the axis by its top 1 word only.
E.3.2 Qualitative observation
In Table 1, we observed examples of consecutive axes extracted from the Axis Tour embeddings with $k=100$. These include the 27th axis (the top five words are france, french, le, paris, and du, so the France axis), the 105th axis (the top five words are lawsuits, lawsuit, litigation, suits, and suit, so the lawsuit axis), and the 241st axis (the top five words are piano, violin, cello, percussion, and orchestral, so the music instruments axis). Table 7 shows the axes close to these axes for $k=1,10,1000$. Note that the axis indices change depending on the results of each Axis Tour.
For each $k$, even the same axis shows significant differences in the nearby axes. For example, at $k=1$, in the bottom row, the meaning of the 227th axis is female, but since the top 1 word is actress, the axis is adjacent to the axes whose meanings are award and movie. This shows the disadvantage of the Axis Tour with $k=1$. At $k=10$, the top row shows that there are more axes related to personal names than to countries near the France axis compared to $k=100$. At $k=1000$, it is interesting to see that the axes France and musical instruments, which are far apart at $k=100$, are close together.
Note that there is a selection bias in this comparison, as we use understandable examples for $k=100$ in Table 1 to compare with $k=1,10,1000$.
Figure 11: The performance of dimensionality reduction for the embeddings including Skewness Sort Projection and Random Order Projection with $\alpha=1/3$. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks.
E.3.3 Dimensionality reduction
Figure 10 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the PCA-transformed embeddings and the Axis Tour embeddings with $k=1,10,100,1000$ and $\alpha=1/3$.
From Fig. 10 we can see that the performance of the Axis Tour embeddings changes for each task, depending on $k$. For example, when comparing across all $k$, $k=1000$ shows good performance on categorization tasks and poor performance on word similarity tasks. In contrast, in lower dimensions, $k=1$ performs better on word similarity tasks than $k=1000$, but shows worse performance on categorization tasks. While $k=100$ does not always show the top performance in all three tasks, it consistently shows stable performance.
These results suggest that the quality of low-dimensional embeddings by the Axis Tour embeddings depends on $k$ for axis embedding.
E.4 Dimensionality reduction by projection for Skewness Sort and Rand Order
The dimensionality reduction by projection in Section 4.3 can be applied not only to Axis Tour, but also to Skewness Sort and Random Order. Therefore, this section compares the dimensionality reduction method with those used for Skewness Sort and Random Order in Section 5.3, where the axes are selected sequentially from the first to last.
For the sake of explanation, let the embedding matrices for Skewness Sort and Random Order be denoted as $\mathbf{S}_{\text{skew}}$ and $\mathbf{S}_{\text{rand}}$, respectively. Let the skewness of the $\ell$-th axis for each be $\gamma_{\text{skew},\ell},\gamma_{\text{rand},\ell}\in\mathbb{R}$. By the definition of Skewness Sort, $\gamma_{\text{skew},\ell}\geq 0$. Similar to Section 4.3, we consider reducing the dimensions from $d$ to $p(\leq d)$, divide $[d]$ into $p$ equal-length intervals, and use the index set for the $r$-th interval $I_{r}=\{a_{r},\ldots,b_{r}\}\,(a_{r},b_{r}\in[d],a_{r}\leq b_{r})$.
E.4.1 Skewness Sort Projection
Given $\gamma_{\text{skew},\ell}\geq 0$, similar to $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ in (5), we define a unit vector $\mathbf{f}_{\text{skew},r}=(f_{\text{skew},r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}_{\geq 0}^{d}$ for each $I_{r}$ as follows:
$f_{\text{skew},r}^{(\ell)}=\begin{cases}\frac{\gamma_{\text{skew},\ell}^{\alpha}}{\sqrt{\sum_{m=a_{r}}^{b_{r}}\gamma_{\text{skew},m}^{2\alpha}}}&\text{for }\ell\in I_{r}\\ 0&\text{otherwise},\end{cases}$ (23)
where $\alpha\in\mathbb{R}_{\geq 0}$. We then define $\mathbf{F}_{\text{skew}}:=[\mathbf{f}_{\text{skew},1},\ldots,\mathbf{f}_{\text{skew},p}]\in\mathbb{R}^{d\times p}$ and obtain the $p$-dimensional embedding matrix $\mathbf{S}_{\text{skew}}\mathbf{F}_{\text{skew}}\in\mathbb{R}^{n\times p}$. We call these embeddings as Skewness Sort Projection.
E.4.2 Random Order Projection
For the Random Order embeddings, some axes may have $\gamma_{\text{rand},\ell}<0$. In this case, $\gamma_{\text{skew},\ell}^{\alpha}$ may not be defined in $\mathbb{R}$ for some $\alpha\in\mathbb{R}_{\geq 0}$. Consequently, we cannot directly use the definition of $\mathbf{f}_{r}$ as a vector for projection. Therefore, even for the Random Order embedding matrix, we flip the sign of each axis to ensure that the skewness is positive, thereby defining a matrix for projection [^17] If the skewness of all axes in the Random Order matrix were initially set to be positive, this complex procedure could be avoided. However, our setup assumes that the sign and order of the axes are arbitrary after the ICA transformation, so the skewness of each axis of the matrix is not positive by default.. To do this, we define a unit vector $\mathbf{f}_{\text{rand},r}=(f_{\text{rand},r}^{(\ell)})_{\ell=1}^{d}\in\mathbb{R}^{d}$ for each interval $I_{r}$ as follows:
$f_{\text{rand},r}^{(\ell)}=\begin{cases}\frac{\text{sgn}(\gamma_{\text{rand},\ell})|\gamma_{\text{rand},\ell}|^{\alpha}}{\sqrt{\sum_{m=a_{r}}^{b_{r}}\gamma_{\text{rand},m}^{2\alpha}}}&\text{for }\ell\in I_{r}\\ 0&\text{otherwise},\end{cases}$ (24)
where $\alpha\in\mathbb{R}_{\geq 0}$ and $\text{sgn}(\cdot)$ represents the sign function. We then define $\mathbf{F}_{\text{rand}}:=[\mathbf{f}_{\text{rand},1},\ldots,\mathbf{f}_{\text{rand},p}]\in\mathbb{R}^{d\times p}$ and obtain the $p$-dimensional embedding matrix $\mathbf{S}_{\text{rand}}\mathbf{F}_{\text{rand}}\in\mathbb{R}^{n\times p}$. We call these embeddings as Random Order Projection.
E.4.3 Results
Similar to Axis Tour, we set the default value of $\alpha$ to $\alpha=1/3$ for both Skewness Sort Projection and Random Order Projection. Figure 11 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the embeddings.
Dimensionality reduction for Axis Tour is still better than or nearly equivalent to Skewness Sort Projection and Random Order Projection for most dimensions in each task. Similar to the results in Section 5.3, these results also suggest the efficiency of dimensionality reduction for Axis Tour.
Note that both Skewness Sort Projection and Random Order Projection show performance improvements over Skewness Sort and Random Order in most dimensions. In particular, both show even better performance than PCA on word similarity tasks. This suggests that they are stronger baseline methods.
Interestingly, despite the lower performance of Random Order in analogy tasks, Random Order Projection performs slightly better than Axis Tour and PCA at $p=50,100$. Considering that Axis Tour showed only equivalent performance to PCA for even different $\alpha$ in Fig. 8 or $k$ in Fig. 10, this result suggests that by applying clustering or similar techniques to axis embeddings, we may obtain more effective low-dimensional vectors.
E.4.4 Setting of $p=d\,(=300)$
When $p=d$, similar to Skewness Sort and Random Order, the performance of Skewness Sort Projection and Random Order Projection for each task is equivalent to that of PCA.
For Skewness Sort Projection, from (23) and the discussion in Appendix B.1, it can be shown that when $p=d$, we have $\mathbf{F}_{\text{skew}}=\mathbf{I}$, and therefore $\mathbf{S}_{\text{skew}}\mathbf{F}_{\text{skew}}=\mathbf{S}_{\text{skew}}\mathbf{I}=\mathbf{S}_{\text{skew}}$.
In preparation for Random Order Projection, we denote $\mathbf{S}_{\text{rand},\geq 0}$ as the matrix obtained by flipping the sign of each axis in $\mathbf{S}_{\text{rand}}$ so that the skewness is positive. It then follows from the discussion in Appendix E.4.2 that, similar to the Skewness Sort Projection, $\mathbf{S}_{\text{rand}}\mathbf{F}_{\text{rand}}=\mathbf{S}_{\text{rand},\geq 0}\mathbf{I}=\mathbf{S}_{\text{rand},\geq 0}$ when $p=d$. Thus, we can see that the performance of $\mathbf{S}_{\text{rand},\geq 0}$ is equivalent to that of $\mathbf{S}_{\text{rand}}$ since $\mathbf{S}_{\text{rand},\geq 0}$ is derived from $\mathbf{S}_{\text{rand}}$ by applying the orthogonal matrix to flip the sign of each axis.
     50 51 52 53 54 55 56 57 58 59
     frustration mindset aspects problems deteriorating rise went was sees given
     anger philosophy factors difficulties evolving rising slipped has gets receive
     resentment attitudes variables hardships urgent spike stumbled requires embraces received
     indignation attitude facets shortcomings rapidly soar ran had creates give
     unhappiness worldview elements troubles dire rises drifted is inspires receiving
     100 101 102 103 104 105 106 107 108 109
     gameplay gambling filly livestock planting birds dog babies father people
     multiplayer casino colt cattle seedlings species dogs infants mother americans
     capcom casinos ##/##-mile farm blooms fish puppy newborn siblings britons
     gamers blackjack mare dairy vines bird pet newborns son patrons
     zelda gamblers gr.3 cows shrub turtles cats baby grandparents viewers
     150 151 152 153 154 155 156 157 158 159
     dh’ing pitched assists nba pointer header everton cambridgeshire milford allegheny
     batting.### #/#-inning steals rondo timeout footed arsenal hertfordshire westerly harrisburg
     ss/2b innings points pistons foul deflected tottenham oxfordshire marlborough pennsylvania
     of/1b ##/#innings .###_##-#-## nets jumper angled fulham staffordshire bridgewater pa.
     2b/of outing rebounds nuggets halftime keeper anfield buckinghamshire amherst penn
     200 201 202 203 204 205 206 207 208 209
     proteins psn### pills beer restaurant sauce diet instructor school sociology
     protein map#### drugs whiskey grill soup diets teach teachers anthropology
     enzyme ap##### heroin lager bistro pasta dieting instruction teacher undergraduate
     molecule gw###### morphine beers diners sauces nutrition beginners elementary postgraduate
     intracellular ly####### ecstasy brewery restaurants meatballs carbohydrate training pupils phd
     250 251 252 253 254 255 256 257 258 259
     workers gdp budget payments tax senate primaries bush mps ontario
     #.##/hour economists cuts payment taxes bill democratic obama labour bc
     ##.##/hr inflation deficit pay surcharge d-tex. voters clinton commons saskatchewan
     9/hour cpi budgets paycheck gst rep. election cheney tory alberta
     wages ism austerity deductible levies r-md. republican putin tories canada
(a) word2vec
     50 51 52 53 54 55 56 57 58 59
     the_1336 air_7 rail_0 of_704 emerge_2 face_2 make_21 see_9 truth_0 takes_1
     the_1335 aviation_0 service_0 the_900 to_350 emerge_0 look_2 know_3 truth_5 a_202
     ’_306 air_5 express_0 the_1609 a_281 find_1 double_5 knew_0 truth_2 maps_0
     s_242 plane_1 the_191 the_901 the_723 yourself_0 sure_2 aware_5 share_10 sweep_0
     it_137 jet_0 trans_0 the_1159 turn_3 to_189 make_26 know_5 truth_1 chronic_0
     100 101 102 103 104 105 106 107 108 109
     liverpool_1 lineup_1 some_8 little_2 five_0 four_6 2_9 both_1 between_12 agreement_0
     annum_1 group_16 some_43 bit_2 two_14 six_0 two_47 both_10 between_9 an_82
     shower_0 band_1 some_21 few_8 two_25 four_15 2_1 both_9 relations_1 struck_1
     ,_755 group_2 many_5 bit_1 few_12 five_1 second_11 two_36 between_10 reached_1
     liverpool_2 number_10 some_27 bit_0 three_9 five_6 two_0 both_6 bonding_0 talks_2
     150 151 152 153 154 155 156 157 158 159
     ,_115 ,_432 ,_1022 ._1512 ._114 a_159 r_0 asher_0 michael_1 emma_0
     ,_701 ,_901 ,_181 ._895 ._1525 r_1 d_12 hilton_0 james_0 cindy_0
     ,_172 ,_876 ,_180 ._1260 ._110 h_1 p_8 colt_0 peter_1 mare_0
     ,_1526 ,_1561 ,_408 ._1339 ._1575 ._381 m_12 ##nie_4 mike_3 kate_0
     ,_1527 ,_1367 ,_1024 ._1055 ._492 ._600 v_0 jesse_0 mike_1 beth_0
     200 201 202 203 204 205 206 207 208 209
     software_1 services_3 limited_1 least_0 cannot_2 also_53 schools_5 bachelor_0 workers_4 users_2
     software_2 capabilities_0 the_271 must_2 a_276 -_613 students_4 degree_0 jobs_2 investors_7
     hardware_0 services_2 are_26 a_98 any_10 cents_2 schools_1 the_739 workers_1 investors_1
     software_0 services_4 to_141 [SEP]_190 in_256 government_29 student_2 major_8 workers_2 users_0
     free_3 services_7 a_116 of_116 to_300 20_14 school_2 subjects_0 workers_10 users_3
     250 251 252 253 254 255 256 257 258 259
     time_9 year_57 hour_3 year_51 day_14 thursday_14 month_10 european_1 world_27 ##8_3
     based_5 year_68 weeks_3 year_37 day_8 tuesday_6 week_10 european_6 world_31 their_17
     based_4 year_56 minutes_1 year_15 day_7 monday_8 year_67 european_0 world_7 nations_0
     lined_2 year_59 hours_4 year_72 day_20 wednesday_3 week_0 european_3 world_17 countries_0
     level_0 year_69 hour_5 year_13 day_2 thursday_0 saturday_0 european_5 world_5 [SEP]_249
(b) BERT
Table 8: Semantic continuity of axes by Axis Tour for normalized ICA-transformed embeddings. We apply Axis Tour to 300-dimensional word2vec and 768-dimensional BERT. For the 50th, 100th, 150th, 200th, and 250th axes, we extract ten consecutive axes from each of these axes and display the top five words for each of the extracted axes.
(a) The 50th axis to 59th axis
(b) The 150th axis to 159th axis
(c) The 250th axis to 259th axis
Figure 12: For word2vec, the scatterplots of the two-dimensional projections for the axes from the 50th to the 59th, from the 150th to the 159th, and from the 250th to the 259th in Table 8(a).
(a) The 50th axis to 59th axis
(b) The 150th axis to 159th axis
(c) The 250th axis to 259th axis
Figure 13: For BERT, the scatterplots of the two-dimensional projections for the axes from the 50th to the 59th, from the 150th to the 159th, and from the 250th to the 259th in Table 8(b).
Axis Tour Skewness Sort
Fig. $d_{I}$ $c_{I}$ $d_{I}$ $c_{I}$
12 0.70 0.30 0.55 0.00
12 0.79 0.34 0.68 0.13
12 0.80 0.27 0.71 0.05
13 0.60 0.28 0.59 0.11
13 0.71 0.33 0.60 0.06
13 0.64 0.31 0.53 0.07
Table 9: The values of $d_{I}$ and $c_{I}$ for Figs. 12 and 13.
E.5 Other embeddings
In Section 5, we used 300-dimensional GloVe [^18] The embeddings can be downloaded here: https://nlp.stanford.edu/data/glove.6B.zip. with $n=400{,}000$. In this section, we extend our experiments to other embeddings. We used 300-dimensional word2vec [^19] https://code.google.com/archive/p/word2vec/ Mikolov et al. (2013b) as static embeddings and 768-dimensional BERT [^20] https://huggingface.co/bert-base-uncased Devlin et al. (2019) from the Hugging Face transformers library Wolf et al. (2020) as dynamic embeddings.
For word2vec, given the original vocabulary size of three million, we selected only the top $40{,}000$ words based on frequency after converting all words to lowercase. The word frequency information was obtained using wordfreq Speer (2022).
For BERT, we first input sentences from the One Billion Word Benchmark Chelba et al. (2014) into BERT, and then used the first $40{,}000$ tokens, including special tokens like [CLS] and [SEP]. It is important to note that the embeddings are different even for identical tokens, so we differentiated tokens like cat as cat_0, cat_1, and so on.
Similar to GloVe, for both word2vec and BERT, we set $k=100$ as the hyperparameter of the axis embedding for Axis Tour.
E.5.1 Qualitative observation
Tables 8(a) and 8(b) show the examples of the axes of the Axis Tour embeddings for word2vec and BERT, respectively. We can also see the semantic continuity of the axes of the Axis Tour embeddings for word2vec in Table 8(a), just as we observed the semantic continuity of those for GloVe in Tables 45 and 6. For BERT, the semantic continuity of the axes is also observed in Table 8(b), although there are axes whose top tokens are the identical tokens. This is due to dynamic embeddings and differs from static embeddings as in GloVe and word2vec.
E.5.2 Scatterplots of Table 8
Figures 12 and 13 show the scatterplots of the two-dimensional projections for the axes of the Axis Tour embeddings for word2vec and BERT in Tables 8(a) and 8(b), respectively. We used the procedure described in Appendix A.1. Similar to Figs. 1 and 6, for both word2vec and BERT, we can see that the top words of the axes are farther from the origin than those of the Skewness Sort, and the meanings of the adjacent axes change continuously.
For Figs. 12 and 13, Table 9 shows the values of $d_{I}$ and $c_{I}$ as the evaluation metrics defined in Appendix A.3. Similar to Table 2, for both word2vec and BERT, Axis Tour shows higher values for both $d_{I}$ and $c_{I}$ than Skewness Sort, indicating better projection quality.
E.5.3 Cosine similarity between adjacent axis embeddings
(a) word2vec
(b) BERT
Figure 14: Histograms of $\cos{(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})}$ for word2vec and BERT. For BERT, we sampled 768 random words and drew the normal distribution $\mathcal{N}(0,1/768)$ instead of 300 words and the normal distribution $\mathcal{N}(0,1/300)$. The rest of the procedure is the same as in Fig. 2.
Figure 14 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ of the Axis Tour embeddings and the baseline embeddings for both word2vec and BERT. Similar to the results for GloVe in Fig. 2, the values of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ are consistently higher in Axis Tour, while this trend is not observed in the other baselines.
Figures 15 and 16 illustrate the relationship between the skewness and the average of consecutive two cosine similarities for Axis Tour and Skewness Sort. Interestingly, while the Axis Tour embeddings for BERT show a strong correlation between the skewness and the average of two cosine similarities in Fig. 16(a), similar to the results for GloVe in Fig. 5(a), the Axis Tour embeddings for word2vec do not show such a correlation in Fig. 15(a).
As future work, it may be interesting to consider the detailed relationship between the skewness and the cosine similarity between adjacent axis embeddings, as well as the relationship between the skewness and the strength of meaning of an axis.
E.5.4 Dimensionality reduction: analogy, word similarity, and categorization tasks
In this section, similar to Section 5.3, we evaluate the performance of dimensionality reduction using the method described in Section 4.3. It is important to note that since BERT embeddings are dynamic embeddings, the token embeddings obtained from sentences in the One Billion Word Benchmark are different from those in each task. Therefore, this evaluation focuses only on word2vec embeddings.
Figure 17 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the embeddings derived from word2vec. Axis Tour outperformed both Random Order and Skewness Sort in each task. In analogy tasks, Axis Tour outperformed PCA in lower dimensions and was nearly equivalent in other dimensions.
Although PCA remains a strong baseline for analogy and categorization tasks because it is an efficient dimensionality reduction method, these results suggest that there is room for performance improvement. However, this study focused primarily on maximizing the semantic continuity of the axes in ICA-transformed embeddings. As mentioned in Limitations, it remains an area for future work to construct more efficient low-dimensional embeddings based on the axes in ICA-transformed embeddings.
Appendix F Topographic ICA (TICA)
Our proposed method, Axis Tour, is related to Topographic Independent Component Analysis (TICA) Kohonen (2001); Hyvärinen et al. (2001a) in terms of ordering the axes of ICA-transformed embeddings, taking into account the sequential relationship between independent components. Therefore, this section first gives a brief overview of TICA. Then, we perform experiments on the TICA embeddings similar to those performed on the Axis Tour embeddings, and then show comparisons of Axis Tour and TICA.
(a) Axis Tour
(b) Skewness Sort
Figure 15: For word2vec, the relationship between the skewness and the average of consecutive two cosines for (a) Axis Tour and (b) Skewness Sort. Spearman’s rank correlation is $-0.04$ for Axis Tour, while it is $0.08$ for Skewness Sort.
(a) Axis Tour
(b) Skewness Sort
Figure 16: For BERT, the relationship between the skewness and the average of consecutive two cosines for (a) Axis Tour and (b) Skewness Sort. Spearman’s rank correlation is $0.50$ for Axis Tour, while it is $-0.10$ for Skewness Sort.
Figure 17: The performance of dimensionality reduction for the embeddings derived from word2vec. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks.
F.1 Overview of TICA
Topographic ICA (TICA) Kohonen (2001); Hyvärinen et al. (2001a) is a variant of linear ICA. While classic ICA assumes independence of the source components [^21] Regarding the indices of the axes, the notation for $\mathbf{f}_{r}=(f_{r}^{(\ell)})_{\ell=1}^{d}$, as seen in the Section 4.3, would be written as $\mathbf{s}=(s^{(\ell)})_{\ell=1}^{d}$. However, for the sake of readability, this section uses the notation $\mathbf{s}=(s_{\ell})_{\ell=1}^{d}$. $\mathbf{s}=(s_{\ell})_{\ell=1}^{d}\in\mathbb{R}^{d}$, TICA allows for positive higher-order correlations $\mathrm{cov}\left(s_{\ell}^{2},s_{m}^{2}\right)$ and assumes that the variances of adjacent sources are correlated. In the probabilistic model of TICA, each variance $\sigma_{\ell}^{2}$ of source component $s_{\ell}$ is generated from the factors $\mathbf{u}=(u_{\ell})_{\ell=1}^{d}\in\mathbb{R}^{d}$ as follows:
$\sigma_{\ell}$ $=\phi\left(\sum_{m=1}^{d}h_{\ell m}u_{m}\right),$ (25)
$s_{\ell}$ $=\sigma_{\ell}z_{\ell},$ (26)
where $\phi$ is some nonlinear function, $h_{\ell m}$ is a symmetric neighborhood relation, and $z_{\ell}$ are mutually independent random variables. To estimate the decomposition matrix $\mathbf{W}=[\mathbf{w}_{1},\ldots,\mathbf{w}_{d}]^{\top}\in\mathbb{R}^{d\times d}$, where $\mathbf{w}_{\ell}\in\mathbb{R}^{d}$, such that $\mathbf{s}=\mathbf{W}\mathbf{x}$ from the mixed signals $\mathbf{x}\in\mathbb{R}^{d}$, TICA maximize the log likelihood
$\log L(\mathbf{W})=\log p\left(\mathbf{s},\mathbf{u}\mid\mathbf{W}\right)=$
$\mathbb{E}_{\mathbf{x}}\left[\log\int\prod_{\ell=1}^{d}\frac{p(z_{\ell})p(u_{\ell})}{\sigma_{\ell}}\left|\det\mathbf{W}\right|d\mathbf{u}\right].$ (27)
By setting $\phi=\sqrt{\cdot}$ and applying some assumptions and approximations, the log likelihood is approximated as
$\log\tilde{L}(\mathbf{W})=\mathbb{E}_{\mathbf{x}}\left[\sum_{m=1}^{d}G\left(\sum_{\ell=1}^{d}h_{\ell m}s_{\ell}^{2}\right)\right]$ (28)
where $G$ is a function approximated by $G(y)=-\beta_{1/2}\sqrt{y}+\beta_{0}$ with constant $\beta_{1/2}=0.8,\beta_{0}=1.2$ in our experiment. For further details on the derivation, see Hyvärinen et al. (2001a).
The optimization of (28) can be done by gradient descent:
$\Delta\mathbf{w}_{\ell}\propto\mathbb{E}_{\mathbf{x}}\left[\mathbf{x}s_{\ell}\sum_{m=1}^{d}h_{\ell m}\frac{dG}{dy}\left(\sum_{m^{\prime}=1}^{d}h_{mm^{\prime}}s_{m^{\prime}}^{2}\right)\right].$ (29)
Note that we have to apply orthonormalization
$\mathbf{W}\leftarrow(\mathbf{WW}^{\top})^{-1/2}\mathbf{W}$ (30)
to the decomposition matrix at each iteration. In accordance with the notation in (1), the extracted components $\mathbf{S}$ are expressed by the obtained $\mathbf{W}$ as
$\mathbf{S}=\mathbf{X}\mathbf{W}^{\top}.$ (31)
In summary, TICA, a kind of ICA, can extract topographic ordered features by using higher-order correlations.
F.2 Experiments
As in classic ICA, we transform the $300$-dimensional GloVe by TICA for our experiments. Following the setup of the 1-D experiment in Hyvärinen et al. (2001a), a neighborhood relation was defined by convolving a rectangular-shaped filter $(\ldots,0,0,0,\overbrace{1,1,\ldots,1,1}^{\mathsf{width}~{}\text{times}},0,0,0,\ldots)$ twice with itself. Hyvärinen et al. (2001a) defined the hyperparameter $\mathsf{width}=5$ for 20 components, so for 300 components, we consider two cases where the width is scaled logarithmically or linearly to $\mathsf{width}=9$ and $\mathsf{width}=75$. These are denoted as TICA9 and TICA75, respectively [^22] For the TICA9-transformed and TICA75-transformed embeddings, similar to the Axis Tour embeddings, we flip the sign of each axis as needed so that the skewness is positive.. We performed $10{,}000$ iterations of gradient descent in both settings.
F.2.1 Qualitative observation
50 51 52 53 54 55 56 57 58 59
coretta david john betjeman prine zuko peter luís de paul
janny paymer rhys-davies john paulk fabricio mcwhinney joao jorge juan
duncum pannick naber blystone kerrys farriss katis monteiro enrique pablo
strouse costabile banville bucchino spinello partida lindroos pereira josé mendoza
livengood zdrilic koepp newcombe pessoa rionda fatialofa joão juan scharner
100 101 102 103 104 105 106 107 108 109
sompolno lb4 : nehzatabad anobiidae tburr poeciliidae onoba wasbir xerocomus
drużbice 3.7995 http://www.***.org poshtkuh ethmia ***.com dipodidae sinezona gurubacharya firered
wartkowice 4-97 http://www.***.com eqbal-e grallariidae ***@***.com pipridae anatoma tolkun quilodran
dzierzgowo 13/11 http://www.***.com zahray-ye sclerosomatidae ***@***.com anthidium cantharidus aomar nishadham
zanjanrud , www.***.com khatunabad leiothrichidae ***.com dendrobatidae syrnola by kirka
150 151 152 153 154 155 156 157 158 159
rilla singen bijar hanus mamat boyolali dammit hockeyallsvenskan v-league
vard referees_mike lali goldring ubayy myogenic gono swac 94-86 lausanne-sport
eguchi jennys prete leura kazagham ragnvald hines ritholtz gøta nasl
kyeong regazzoni markazi larmer munnetra yamba strayhorn yorick 86-82 whl
achiou x13 garone similis terefe swabs sabzevar elmyra second-tier divisione
200 201 202 203 204 205 206 207 208 209
two-week 35-year epigraphic 2001-2003 snags sponsorship spectator frantic appeals translations
six-week 50-year fogies 1998-2001 objections logo bhb hysterical frivolous translation
a-half 60-year old 1996-1998 disagreements pitfall moot liquidator economized pashto
five 25-year blighty 1999-2002 procedural tagline unfit quixotic appeal word
six 40-year gutnish 1995-1997 weaknesses drawcard sjc counting 360-degree arabic
250 251 252 253 254 255 256 257 258 259
ltte gam czech russian anatoly korean indian serb romania minh
ira aceh poland moscow ossetia south mukherjee croatia moldova nsanje
eelam milf krakow vasily ossetian jang bengal zagreb inyo chennai
tamil rebel warsaw sergei oleg korea jharkhand serbian iceland srinivasan
fein mnlf polish vladimir interfax kim bihar croatian norway chikwawa
(a) TICA9
50 51 52 53 54 55 56 57 58 59
gartenberg faden trescothick .652 billcliff dani lentulus s.s.d. dakhlallah matthewson
nart tilles bopara ramnaresh plzen siggins sivasspor torrin ollivier feiner
radziwill kolderie harmse razaq hakohen feith najdi sastra clemetson hatchell
stachowski strada briones taufeeq ashraful daugaard epps vendémiaire herrin carrickmacross
recio durrant ruvo sarwan xiangfan chedid sohlman wenping sampley lynskey
100 101 102 103 104 105 106 107 108 109
vnccpbn lomartire kihn jędrzejów cooperacion ***-***-**** prusice extranjero :53 sa/b
ustbimp rsst1 apenas pleszew cvik ohernandez rmartinez finalmente qe5 ***.com
gromada 36.41 zgray sokołów hrvatske .367 kmorales ***@***.com jgreig ***@***.com
a_21 darreh-ye divergens przysucha 0255 (928) alifereti cantidad fedbud ***@***.com
**-***-**** immatures nohv raciąż entidad cjones separado ***.net rxe5 ***@***.com
150 151 152 153 154 155 156 157 158 159
reteamed leaf-sheaths risen auguring i backward-compatible languishes thursday.the expeditiously surged
unbeknown upsmanship surged two-song n’t satisfactorily 5.4 one-disc resolved soared
unbeknownst udalls obrija 18-yard ’ve outwardly 5.8 50-50 sometime leapt
co-operate great-grandchild soared unproduced happen co-existed 4.9 32-page commence leaped
re-connect nineteen 07:20 epithalamium presume satisifed 4.1 hefty be leaps
200 201 202 203 204 205 206 207 208 209
expert ceremony exams vandalism adjacent rocking hazards recommendations 27th cd-rom
experts celebrating mathematics campus blocks bed investigators advisory 13th mods
feasibility eve ribbon antisocial building sitting detect 301st 10th evolution
forensic parade math graffiti erected padded exposures advancement 15th famines
study commemorating school streetscape built pounding remediation psc 23rd experimental
250 251 252 253 254 255 256 257 258 259
titanic saratoga smolensk ms-dos 2g pol supermarine ***.com ump under-17
clot 1772 hasselblad pelts enos muse xk120 aretha iwf huon
vessel 1793 bonifacio production-based blandest plasterboard ground-attack banana icing pixie
tanker thoroughbred cornplanter conquering pagan spray turbocharged canes mapo southwood
stotesbury 1794 1314 1405 istar bmw madelyn amazon vedanta trikke
(b) TICA75
Table 10: Semantic continuity of axes for normalized TICA9-transformed and TICA75-transformed embeddings. For the 50th, 100th, 150th, 200th, and 250th axes, we extract ten consecutive axes from each of these axes and display the top five words for each of the extracted axes.
(a) The 50th axis to the 59th axis
(b) The 150th axis to the 159th axis
(c) The 250th axis to the 259th axis
Figure 18: For TICA9, the scatterplots of the two-dimensional projections for the axes from the 50th to the 59th, from the 150th to the 159th, and from the 250th to the 259th in Table 10(a).
(a) The 50th axis to the 59th axis
(b) The 150th axis to the 159th axis
(c) The 250th axis to the 259th axis
Figure 19: For TICA75, the scatterplots of the two-dimensional projections for the axes from the 50th to the 59th, from the 150th to the 159th, and from the 250th to the 259th in Table 10(b).
(a) The 50th axis to the 59th axis
(b) The 150th axis to the 159th axis
(c) The 250th axis to the 259th axis
Figure 20: For Axis Tour, the scatterplots of the two-dimensional projections for the axes from the 50th to the 59th in Table 4, from the 150th to the 159th in Table 5, and from the 250th to the 259th in Table 6.
Table 10 shows the examples of the axes of the TICA9-transformed embeddings and the TICA75-transformed embeddings. Similar to Axis Tour, we observe the semantic continuity of the axes in some subintervals. For example, the 250th to the 259th axes in Table 10(a) show the semantic continuity with respect to countries. Moreover, for these examples in Table 10, TICA9 seems to show the clearer semantic continuity of the axes compared to TICA75. Note that URLs, email addresses and phone numbers are anonymized in Table 10.
F.2.2 Scatterplots of Table 10
Figures 18 and 19 show the scatterplots of the two-dimensional projections for the the axes of the TICA9-transformed embeddings and the TICA75-transformed embeddings in Tables 10(a) and 10(b). In these examples, for both TICA9 and TICA75, the interpretability of the embedding space from the scatterplots decreases as the top words of the axes show significant movement or approach the origin. This property is particularly noticeable in TICA75. For example, in Fig. 19, the top words of the 159th axis are far from the axis: surged and soared are concentrated near the 156th axis, and leapt and leaped are concentrated near the origin.
Fig. embeddings $d_{I}$ Skew. $d_{I}$ diff
18 TICA9 0.61 0.59 0.02
18 0.33 0.35 -0.02
18 0.57 0.54 0.03
19 TICA75 0.35 0.31 0.04
19 0.44 0.44 0.00
19 0.31 0.34 -0.03
20 Axis Tour 0.78 0.73 0.05
20 0.72 0.58 0.14
20 0.57 0.52 0.05
Avg. TICA9 0.59 0.54 0.05
TICA75 0.41 0.41 0.00
Axis Tour 0.68 0.59 0.09
Table 11: The values of $d_{I}$ for Figs. 1819 and 20, and the average values of $d_{I}$ over 300 subintervals with $|I|=10$. Skew. stands for Skewness Sort.
Fig. embeddings $c_{I}$ Skew. $c_{I}$ diff
18 TICA9 0.35 0.21 0.14
18 0.09 0.03 0.06
18 0.25 0.16 0.09
19 TICA75 0.07 0.06 0.01
19 0.06 0.11 -0.05
19 0.09 0.06 0.03
20 Axis Tour 0.27 0.11 0.16
20 0.27 0.07 0.20
20 0.12 0.04 0.08
Avg. TICA9 0.23 0.14 0.09
TICA75 0.13 0.13 0.00
Axis Tour 0.24 0.09 0.15
Table 12: The values of $c_{I}$ for Figs. 1819 and 20, and the average values of $c_{I}$ over 300 subintervals with $|I|=10$. Skew. stands for Skewness Sort.
Figure 21: Histograms of cosine similarities between adjacent axis embeddings for the TICA-transformed embeddings, the Axis Tour embeddings and the Random Order embeddings.
Figure 22: Scatterplots of $\cos{(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})}$ for Random Order, Axis Tour, TICA9 and TICA75.
Figure 23: The performance of dimensionality reduction for the PCA-transformed embeddings, the Axis Tour embeddings with $\alpha=1/3$, the TICA-transformed embeddings, and the TICA Projection embeddings with $\alpha=1/3$. Each value represents the average of 30 analogy tasks, 8 word similarity tasks, or 6 categorization tasks.
Figure 24: Histograms of higher-order correlations between adjacent axes for the TICA-transformed embeddings, the Axis Tour embeddings, and the Random Order embeddings.
F.3 Comparisons of Axis Tour and TICA
F.3.1 Scatterplots
This section compares the scatterplots of the TICA9-transformed embeddings in Fig. 18 and the TICA75-transformed embeddings in Fig. 19 with those of Axis Tour. Since TICA9, TICA75, and Axis Tour yield different embeddings, we also compare them based on Skewness Sort.
First, similar to Figs. 18 and 19, we use 300-dimensional GloVe and show the scatterplots for comparison232323 Note that Figs. 1 and  6 show the scatterplots for the illustrative examples, so it is unfair to compare them directly with those of TICA. in Fig. 20, which shows the two-dimensional projections for the axes of the Axis Tour embeddings.
Then Table 11 shows the values of $d_{I}$ defined in Appendix A.3 for Figs. 1819, and 20, and the average values of $d_{I}$ over 300 subintervals with $|I|=10$. In these examples, Axis Tour shows larger $d_{I}$ values and larger differences with Skewness Sort compared to TICA9 and TICA75. These results are also observed for the average values. In addition, we can see the cases where $d_{I}$ is smaller than that of Skewness Sort in both TICA9 and TICA75, such as Figs. 18 and 19.
Table 12 shows the values of $c_{I}$ defined242424 For the TICA9- and TICA75-transformed embeddings, as with Axis Tour embeddings, $c_{I}$ can be computed by the cosine similarities of adjacent axis embeddings. in Appendix A.3 for Figs. 1819, and 20, and the average values of $c_{I}$ over 300 subintervals with $|I|=10$. In these examples, TICA9 shows the $c_{I}$ values comparable to Axis Tour. In fact, their average values are almost the same. However, TICA9 shows larger $c_{I}$ values even when Skewness Sort is performed on $I$. This suggests that for the TICA-transformed embeddings, the axes with similar meanings tend to cluster locally rather than being ordered so that the meanings are continuous. This hypothesis could explain why, in the scatterplots such as Figs. 18 and 19, the top words of the axes show significant movement or approach the origin.
In TICA75, the $c_{I}$ values are lower, and the difference in the average value compared to Skewness Sort is 0. This indicates that the axes with less similar meanings are grouped together in $I$, resulting in no difference in $c_{I}$ between TICA and Skewness Sort.
F.3.2 Cosine similarity and higher-order correlation
Figure 21 shows the histograms of $\cos(\mathbf{v}_{\ell},\mathbf{v}_{\ell+1})$ of the TICA-transformed embeddings, the Axis Tour embeddings and the Rnadom Order embeddings. Similar to Axis Tour, The distribution for TICA shifts towards a more positive mean than that of Random Order. Note that while TICA9 has an average close to that of Axis Tour, it has a significantly higher variance of cosine similarity.
To highlight this, Fig. 22 shows scatterplots of cosine similarities between adjacent axis embeddings. The variance is $0.05$ for Random Order, $0.12$ for Axis Tour, $0.25$ for TICA9, and 0.15 for TICA75. The variance for TICA9 is significantly larger than that for Axis Tour, which indicates that the semantic continuity of the axes changes more drastically in the TICA9-transformed embeddings.
Figure 24 shows the histograms of higher-order correlations between adjacent axes of the TICA-transformed embeddings, the Axis Tour embeddings and the Rnadom Order embeddings. We can see that the average higher-order correlation for TICA is higher than that for Axis Tour, reflecting the learning settings of TICA.
F.3.3 Dimensionality reduction: analogy, word similarity, and categorization tasks
This section performs dimensionality reduction for TICA, selecting the axes sequentially starting from the first, similar to Random Order and Skewness Sort.
As we saw in Appendix E.4, we defined Random Order Projection and Skewness Sort Projection by the dimensionality reduction process similar to that of Axis Tour. A similar approach can be considered for TICA, and we call this method TICA projection.
Role Prompt
system You are an excellent NLP annotator. Your response should be in JSON format with the key ‘choice’.
user Which of the following words are related to the words [top1 word,$\ldots$, top10 word]. Answer A or B.
A. [Axis Tour top1 word,$\ldots$, Axis Tour top10 word]
B. [Skewness Sort top1 word,$\ldots$, Skewness Sort top10 word]
Table 13: Prompts for the GPT models. Since the Axis Tour embeddings and the Skewness Sort embeddings differ solely in the order of the axes, they share $d\,(=300)$ common axes. The top 10 words for each common axis are denoted as [top1 word,$\ldots$, top10 word]. We focus on the next axis of Axis Tour and the next axis of Skewness Sort, based on the common axis. The top 10 words for the Axis Tour axis are denoted as [Axis Tour top1 word,$\ldots$, Axis Tour top10 word], while the top 10 words for the Skewness Sort axis are denoted as [Skewness Sort top1 word,$\ldots$, Skewness Sort top10 word]. We then use the top words for the prompt.
Model Version
GPT-3.5 Turbo gpt-3.5-turbo-0125
GPT-4 Turbo gpt-4-turbo-2024-04-09
GPT-4o gpt-4o-2024-05-13
GPT-4o mini gpt-4o-mini-2024-07-18
Table 14: Version of each GPT model.
Figure 23 shows the average of each task at $p=1,2,5,10,20,50,100,200,300$ for the Axis Tour embeddings, and TICA-transformed embeddings. Axis Tour outperformed both TICA and TICA Projection for most dimensions in each task. Note that similar to the results for Random Order and Skewness Sort in Appendix E.4, TICA Projection shows performance improvements over TICA in most dimensions. These results demonstrate the utility of projection-based dimensionality reduction even in TICA, an algorithm that relaxes the assumption of statistical indepence for ICA.
Appendix G Details of the quantitative evaluation of semantic continuity by GPT models
This section provides details on the GPT models, accessed via the OpenAI API, and the prompts used in Section 5.2.2.
The versions of the GPT models used in the experiments are listed in Table 14. The prompts used for the experiments are shown in Table 13. In Fig. 3, the responses obtained from each model using these prompts are aggregated and shown. Note that GPT-4 Turbo failed to provide responses relevant to either Axis Tour or Skewness Sort for two queries.